### v1

# On the Verification of Control Flow Attestation Evidence

Abstract
Remote run-time attestation methods, including Control Flow Attestation (CFA) and Data Flow Attestation (DFA), have been proposed to generate precise evidence of execution’s control flow path (in CFA) and optionally execution data inputs (in DFA) on a remote and potentially compromised embedded device, hereby referred to as a Prover ($\sf{\mathcal{P}rv}$). Recent advances in run-time attestation architectures are also able to guarantee that a remote Verifier ($\sf{\mathcal{V}rf}$) reliably receives this evidence from $\sf{\mathcal{P}rv}$, even when $\sf{\mathcal{P}rv}$’s software state is fully compromised. This, in theory, enables secure “run-time auditing” in addition to best-effort attestation, i.e., it guarantees that $\sf{\mathcal{V}rf}$ can examine execution evidence to identify previously unknown compromises as soon as they are exploited, pinpoint their root cause(s), and remediate them. However, prior work has for the most part focused on securely implementing $\sf{\mathcal{P}rv}$’s root of trust (responsible for generating authentic run-time evidence), leaving $\sf{\mathcal{V}rf}$’s perspective in this security service unexplored.
In this work, we argue that run-time attestation and auditing are only truly useful if $\sf{\mathcal{V}rf}$ can effectively analyze received evidence. From this premise, we characterize different types of evidence produced by existing run-time attestation/auditing architectures in terms of $\sf{\mathcal{V}rf}$’s ability to detect and remediate (previously unknown) vulnerabilities. As a case study for practical uses of run-time evidence by $\sf{\mathcal{V}rf}$, we propose SABRE: a Security Analysis and Binary Repair Engine. SABRE showcases how $\sf{\mathcal{V}rf}$ can systematically leverage run-time evidence to detect control flow attacks, pinpoint corrupted control data and specific instructions used to corrupt them, and leverage this evidence to automatically generate binary patches to buffer overflow and use-after-free vulnerabilities without source code knowledge.
Index Terms:
Control Flow Attestation, Systems Security, Embedded Systems
1 Introduction
Embedded devices exist at the edge of larger systems, often serving as remotely operated sensors and on-demand actuators. Due to low energy and cost requirements, they are commonly implemented using low-end micro-controller units (MCUs) that lack the architectural security features of application computers (e.g., virtual memory or memory management units). Consequently, MCUs are relatively more vulnerable and frequently targeted by cyber attacks [1, 2].
Given their low budget for preventative security features, MCUs need inexpensive means to prove their software integrity to back-ends that rely on their services (e.g., a device owner or control center). Towards this goal, Remote Attestation (RA[3] has been proposed to give a resource-rich back-end – called a Verifier ($\sf{\mathcal{V}rf}$) – the ability to remotely assess the software state of a resource-constrained Prover MCU ($\sf{\mathcal{P}rv}$). In RA, $\sf{\mathcal{V}rf}$ challenges $\sf{\mathcal{P}rv}$ to return a cryptographic proof of $\sf{\mathcal{P}rv}$’s currently installed software binary. Only after receiving an authentic proof will $\sf{\mathcal{V}rf}$ trust that $\sf{\mathcal{P}rv}$ is installed with the correct software.
Although RA can effectively detect illegal code modifications, it cannot detect run-time attacks that do not modify code [4]. For example, an adversary ($\sf{\mathcal{A}dv}$) could hijack the control flow of a program by exploiting a memory vulnerability (e.g., a buffer overflow [5]) to corrupt control data, including return addresses, indirect jump targets, and function pointers. Consequently, $\sf{\mathcal{A}dv}$ can skip access controls checks [6] or chain the out-of-order execution of sub-sequences of instructions to implement (often Turing-complete) malicious behavior [7, 8]. Since these attacks do not modify code, they are oblivious to RA.
To address this limitation, RA can be extended into Control Flow Attestation (CFA). In CFA, $\sf{\mathcal{V}rf}$ obtains cryptographic proof of both the installed software image and the exact control flow path followed in the most recent execution of the software [9]. This evidence and the respective proof are generated/authenticated by a root of trust (RoT) within $\sf{\mathcal{P}rv}$ that securely records control flow transfer destinations. Therefore, CFA evidence can, in principle, allow $\sf{\mathcal{V}rf}$ to detect both code modification and out-of-order execution attacks. Recent architectures augment CFA to obtain “Run-time Auditing” [6, 10] by guaranteeing the delivery of CFA evidence to $\sf{\mathcal{V}rf}$ even when $\sf{\mathcal{P}rv}$’s software is compromised. See Sec. 2 for details on CFA and run-time auditing RoTs.
We note, however, that prior work in this area [9] has primarily focused on secure design/implementation of the $\sf{\mathcal{P}rv}$ RoTs, often overlooking $\sf{\mathcal{V}rf}$’s role in run-time attestation/auditing. Current efforts focusing on $\sf{\mathcal{V}rf}$ designs look into hiding control flow paths while enabling simple remote checks via Zero-Knowledge proofs [11] or addressing scalability concerns of building control flow graphs (CFG) for path verification [12].
Yet, despite the promising benefits of CFA, no prior work has systematically analyzed $\sf{\mathcal{V}rf}$ ability to interpret received CFA evidence to identify root cause memory vulnerabilities that lead to attacks and remediate them. These challenges are catalyzed when firmware source codes are unavailable due to containing proprietary dependencies/third-party libraries: an unfortunate common occurrence in embedded system software chains. Remarkably, given the prevalence of memory-unsafe languages in embedded implementations and their various associated exploits and bugs [13], we argue that automated methods to process run-time evidence to identify unknown memory safety vulnerabilities (as soon as they are exploited in control flow attacks) can result in timely detection of otherwise oblivious attacks and elimination of their causes.
While techniques have been proposed for root cause vulnerability analysis in high-end devices (e.g., based on Intel PT traces [14, 15]) there are no frameworks for analysis of execution traces produced by run-time attestation in MCUs. Our work aims to bridge these gaps with the following intended contributions:
  1. (1)
    We analyze and classify the types of run-time evidence generated by existing CFA RoTs. Their trade-offs are examined in detail with particular focus on their ability to support $\sf{\mathcal{V}rf}$’s remote detection and remediation of vulnerabilities.
  2. (2)
    As a case study, we present SABRE: a Security Analysis and Binary Repair Engine based on CFA evidence. SABRE combines binary analysis with CFA run-time evidence to automatically locate and patch buffer overflow and use-after-free vulnerabilities without requiring source-code knowledge. SABRE analysis combines backward definitions analysis along with forward symbolic data-flow analysis to identify such root causes. Based on this analysis and the detected vulnerability, SABRE automatically generates patches and validates their effectiveness.
  3. (3)
    We develop a fully functional and open-source prototype of SABRE with support for binaries of TI MSP430 and ARM Cortex-M MCU families [16]. SABRE prototype is also directly compatible with run-time evidence generated by existing open-source run-time attestation RoTs based on Trusted Execution Environments (TEEs) [10] and custom hardware [6].
2 Background
2.1 Remote Attestation
RA is a challenge-response protocol between two parties: a remote $\sf{\mathcal{V}rf}$ and a potentially compromised $\sf{\mathcal{P}rv}$. It allows $\sf{\mathcal{V}rf}$ to remotely assess $\sf{\mathcal{P}rv}$ trustworthiness by measuring the content of $\sf{\mathcal{P}rv}$ program memory. As depicted in Fig. 1, a typical RA interaction involves the following steps:
  1. (1)
    $\sf{\mathcal{V}rf}$ requests attestation from $\sf{\mathcal{P}rv}$ by sending a cryptographic challenge $Chal$.
  2. (2)
    Upon receiving $Chal$, an RoT in $\sf{\mathcal{P}rv}$ computes an authenticated integrity-ensuring function over its program memory and $Chal$ to produce report $H$.
  3. (3)
    $\sf{\mathcal{P}rv}$ sends the report $H$ back to $\sf{\mathcal{V}rf}$.
  4. (4)
    Upon receiving $H$, $\sf{\mathcal{V}rf}$ checks $H$ against the expected value to determine if $\sf{\mathcal{P}rv}$ has been compromised.
A RoT in $\sf{\mathcal{P}rv}$ performs the authenticated measurement (step 2) by generating an integrity token, e.g., a message authentication code (MAC) or a digital signature, over program memory using a secret key (either pre-shared symmetric key or the private key for a public key $\sf{\mathcal{V}rf}$ knows). Since threat models of RA (and of CFA consequently) assume that $\sf{\mathcal{P}rv}$ is susceptible to full software compromise, the secret key used in this operation must be securely stored by $\sf{\mathcal{P}rv}$’s RoT. Secure storage for the RA secret key implies some level of hardware support for the RoT implementation, ensuring that the key is unmodifiable and inaccessible to any untrusted software running on $\sf{\mathcal{P}rv}$.
Some early RA proposals aimed to generate evidence of $\sf{\mathcal{P}rv}$’s code integrity without hardware support. In this setting, the evidence provided to $\sf{\mathcal{V}rf}$ must be deterministic and have a measurable architectural side-effect (e.g., predictable delivery time [17]). These “software-based” architectures detect attempts to forge RA evidence based on delayed responses from $\sf{\mathcal{P}rv}$. Hence, they require assumptions about deterministic timing characteristics, direct wired interfaces between $\sf{\mathcal{V}rf}$-$\sf{\mathcal{P}rv}$, and predictable latency [17, 18, 19, 20]. Unfortunately, these strict requirements make them incompatible with remote settings [21]. Recent software-based architectures replace the timing requirements for a trusted software boot-strapping phase [22, 23]; additionally, they use software-fault isolation to ensure the integrity of the trusted components at run-time.
Figure 1: RA interaction
Hardware-based RA [24, 25, 26, 27, 28] incorporates dedicated hardware to securely create RA evidence that depends on secrets rather than timing characteristics. This is achieved through external modules (such as custom hardware extensions [27, 28, 25] or TPMs [26]) or RA-dedicated co-processors [24]. While they provide stronger security guarantees than earlier software-based techniques, the added hardware cost can be prohibitive for resource-constrained devices, such as MCUs.
Hybrid RA schemes  [29, 30, 31] use a combination of minimal hardware components and trusted software modules to achieve the security guarantees equivalent to hardware-based RA at a manageable hardware cost. Hybrid approaches typically implement the RA measurement (e.g., MAC or signature functionality) in software, while a low-cost hardware monitor protects the integrity of the RA measurement code and its execution and guarantees the secrecy of the RA cryptographic key(s).
2.2 Run-Time Code Reuse/Control Flow Attacks
[…]
2.2.1 Control Flow Attack Root Causes and their Prevalence in Embedded Systems Software
One class of memory vulnerabilities that can be exploited to launch a control flow attack is a buffer overflow [5]. A buffer overflow occurs when a memory write exceeds the bounds of a buffer. If $\sf{\mathcal{A}dv}$ can cause a buffer overflow and control its input, they can leverage this vulnerability to overwrite adjacent memory that stores indirect branch destinations (e.g., return addresses or function pointers), successfully launching control flow attacks without injecting code directly. See Appendix A.1 for a concrete example of a control flow attack via a buffer overflow.
Buffer overflows are one of the most commonly exploited memory vulnerabilities [36], especially in embedded systems software that is commonly written in unsafe memory languages like C/C++. Some recent examples include an exploited buffer overflow in Realtek chipset’s SDK that allowed remote attackers to launch arbitrary code execution in routers and various IoT devices [37] (consequently resulting in CVE-2021-35395 [38]). Buffer overflow vulnerabilities were also previously disclosed in embedded real-time operating systems (RTOS), leading to arbitrary code execution (e.g., CVE-2020-10023 [39] and CVE-2020-10019 [40] in Zephyr [41]). This style of attack is not limited to return addresses on the stack. For example, each instance of a C++ object includes a function pointer to a “vtable” of the object’s C++ class’s functions [42]. Thus, an $\sf{\mathcal{A}dv}$ could hijack a program’s control flow by causing an overflow into a C++ object’s function pointers. This applies equally to C structs containing function pointers.
Another memory vulnerability that often leads to control flow attacks is use-after-free [43]. In use-after-free, dynamically allocated memory is used after it has been freed, typically leading to a program crash when accidental (i.e., in a non-malicious case). When exploited, it can corrupt control data or launch control flow attacks. For example, when that freed memory previously contained function/vtable pointers (e.g., in the case of dynamically allocated C structs or C++ objects), $\sf{\mathcal{A}dv}$ can modify freed memory so that the reuse of this previously freed memory leads to a control flow hijack. A concrete example of how a use-after-free can lead to control flow hijacking is presented in Appendix A.2.
Use-after-frees were ranked as one of the most dangerous software weaknesses by MITRE [36]. Embedded libraries/firmware that use dynamic allocation [44, 45, 46] can be equally vulnerable use-after-free vulnerabilities [47, 48]. An example of use-after-free leading to arbitrary code execution in embedded platforms includes CVE-2017-14201 [47] that affected Zephyr RTOS [41].
Whether launched by a buffer overflow or a use-after-free vulnerability, such attacks do not require code modifications. Thus, they would remain undetected by classic RA mechanisms.
2.3 Run-time Attestation and Auditing
Run-time Attestation and Auditing are a set of security services that aim to generate and guarantee the delivery of accurate and authentic evidence of a $\sf{\mathcal{P}rv}$’s run-time behavior to a remote $\sf{\mathcal{V}rf}$, thus reflecting to $\sf{\mathcal{V}rf}$ whether any control flow attack has occurred. One type of run-time attestation is Control Flow Attestation (CFA[49, 50, 51, 52, 53, 54, 4, 55, 56, 57, 58], which extends RA to detect run-time attacks in addition to illegal binary modifications. Upon receiving a request to execute some software function along with a RA challenge from $\sf{\mathcal{V}rf}$, $\sf{\mathcal{P}rv}$ executes the requested function, and a RoT in $\sf{\mathcal{P}rv}$ records a trace of this execution into protected memory. In the most fine-grained case, the trace is a control flow log ($\mathcal{C}F_{Log}$) of all control flow transfers that have occurred during the requested execution. After producing $\mathcal{C}F_{Log}$, $\sf{\mathcal{P}rv}$ performs the authenticated RA integrity measurement over the $\mathcal{C}F_{Log}$, the received challenge, and its own program memory to produce a timely and authenticated report. This report attests to the static state of the $\sf{\mathcal{P}rv}$’s program binary and the run-time state within this timely execution. When applicable, data inputs – in data flow attestation (DFA[50, 59, 51] – or produced outputs [60, 61] can also be included in this proof to bind obtained results to the proper execution of respective software functions on expected inputs.
Existing CFA techniques use either (1) binary instrumentation along with TEE support; or (2) custom hardware modifications to generate $\mathcal{C}F_{Log}$ by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that all branch instructions (e.g., jumps, returns, calls, etc.) are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends $\mathcal{C}F_{Log}$ to update the trace with the current branch destination. In hardware-based techniques, custom hardware interfaces with the MCU core to detect branches and records their destinations to a reserved memory region. Recent work [6, 10] has also created mechanisms to implement reliable communication with $\sf{\mathcal{V}rf}$ as a part of the CFA RoT, enabling “run-time auditing” by ensuring evidence delivery even in the presence of malware that infects $\sf{\mathcal{P}rv}$ and refuses to deliver evidence about its actions/presence.
3 Characterizing Run-time Evidence and their Utility
3.1 Evidence Types
The various realizations of CFA RoTs (see Sec. 2) create and manage $\mathcal{C}F_{Log}$ differently, presenting trade-offs between the storage/communication cost and ease of verification/analysis by $\sf{\mathcal{V}rf}$. We classify existing methods in three types: hash-based evidence (E1), verbatim evidence (E2), and hybrid evidence (E3).
Figure 2: Illustration of logging a conditional branch (CB), indirect/direct calls (IC/DC), and returns (RET) in $\mathcal{C}F_{Log}$ with (a) hash-based, (b) verbatim, and (c) hybrid evidence.
3.1.1 Hash-based Evidence (E1)
As control flow transfers occur, branch destination addresses are accumulated into a single hash chain. This is performed by first accumulating the current destination address with the previous hash and then hashing the accumulated result [4, 62, 56, 52], as exemplified in Fig. 2.a. After the first branch occurs, its destination 0xe004 is accumulated into the hash. Since this is the first branch destination, it is accumulated with an initial constant to obtain $H_{0}$. Upon the second branch instruction, its destination address 0xf000 is accumulated with $H_{0}$ to obtain $H_{1}$. This process continues until the end of the requested execution generating a unique value that represents the path. Although simple control flows produce a single hash in E1, programs with more complex control flows (including those with nested loops and branching) might require a set of hash digests [4, 52]. This approach yields minimal storage/transmission overheads (i.e., the fixed size of a hash digest). However, $\sf{\mathcal{V}rf}$ must perform its analysis solely based on the final hash digest and the executable binary.
3.1.2 Verbatim Evidence (E2)
On the other side of the spectrum, more recent CFA techniques advocate for storing and transmitting $\mathcal{C}F_{Log}$ in its entirety [49, 51, 53, 6, 54, 10]. As depicted in Fig. 2.b, the full sequence of destination addresses is maintained in a buffer. Therefore, this approach requires further (lossless) compression or is limited to small and self-contained operations to avoid filling the dedicated storage too quickly. A common strategy is to log simple loops (those without internal branches) with their address once and a count/condition pertaining to the number of iterations performed [6, 49, 54, 10]. Another reduction is obtained by recording statically defined branch destinations as a single bit [51]. Since static destinations can be determined from the binary alone, branch instructions with statically defined destinations can be recorded as a single bit (1/0) to determine if the branch was taken or not. Branches with dynamically defined destinations (e.g., returns and indirect calls) are still recorded with their full addresses. If $\mathcal{C}F_{Log}$ remains larger than available storage despite these optimizations, it can be split into a set of multiple fixed-size slices and streamed to $\sf{\mathcal{V}rf}$, as suggested in [53, 6, 10]. Despite the aforementioned challenges, E2 provides $\sf{\mathcal{V}rf}$ with full control flow path information for analysis.
3.1.3 Hybrid Evidence (E3)
Some techniques implement a combination of E1 and E2. OAT [50] and ARI [63] record forward edges verbatim and accumulate backward edges (returns) into a single hash chain. As shown in Fig. 2.c, the RoT in $\sf{\mathcal{P}rv}$ maintains a trace of forward edges ($T_{F}$) and a hash chain of returns ($H_{R}$). Since 0xf000 is the destination of an indirect call, its full address must be recorded whereas the first destination can be represented by a single bit ‘0’ encoding the ‘not-taken’ conditional branch address. After the attested execution ends, $T_{F}$ and $H_{R}$ are sent to $\sf{\mathcal{V}rf}$.
3.2 $\sf{\mathcal{V}rf}$’s Perspective: Detection & Remediation from E1, E2 & E3
Consider an attack detection strategy in which $\sf{\mathcal{V}rf}$ constructs the CFG of the executable binary, i.e., as suggested in [4, 6, 62, 52, 53, 56, 54]. An initial set of invalid paths is defined as forward edges that do not exist in the statically constructed CFG. As valid indirect targets (e.g., return addresses or indirect calls) are over-approximated using the static CFG [64], $\sf{\mathcal{V}rf}$ may also use $\mathcal{C}F_{Log}$ along with the program’s binary to emulate the execution of the program locally. In this process, a Shadow Stack (SS) can be emulated locally to verify backward edges (return destinations) and data flow analysis can be employed to determine a set of valid call sites for each indirect call, updating the CFG accordingly. We observe feasibility of this verification strategy varies greatly depending on evidence types E1, E2, and E3.
Figure 3: SABRE Workflow
For E1, with the only evidence available being a hash digest, $\sf{\mathcal{V}rf}$ must infer which path has resulted in the digest. To that end, it must generate each possible legal path based on the CFG and use an emulated SS to accumulate the path’s hash during each step of the CFG traversal. However, the complexity of this task grows exponentially with the number of branches, leading to the intractable path explosion problem [65, 66]. Considering anomalous/illegal paths, the analysis is even more challenging. $\sf{\mathcal{V}rf}$ would need to map the hash digest to the exact illegal path. Illegal paths do not exist in the CFG and thus must then be enumerated directly, making it effectively impossible to learn about the exploit behavior from this type of evidence (this precludes any analysis of exploit root causes, as discussed below).
Evidence types E2 and E3 allow $\sf{\mathcal{V}rf}$ to use reported traces to traverse the CFG, checking the reported destination against valid destinations. When there is a mismatch between the reported destination and the set of valid destinations for that branching instruction, $\sf{\mathcal{V}rf}$ determines that $\sf{\mathcal{P}rv}$ executed an invalid path. With both E2 and E3, $\sf{\mathcal{V}rf}$ is able to emulate an SS of return addresses. For verifying the integrity of returns in E2, $\sf{\mathcal{V}rf}$ checks each return destination individually as they are reached in the path traversal. The expected return address is popped from the SS and compared to the return address reported in $\mathcal{C}F_{Log}$, and an invalid path is detected when there is a mismatch. For E3, however, $\sf{\mathcal{V}rf}$ must recompute all returns reached into the final hash chain digest and compare it to the one reported in $\mathcal{C}F_{Log}$. If there is a mismatch of hashes, an invalid path is detected due to one of the return addresses. We note that verification based on E3 is only able to determine that some return address was corrupted. Since actual return addresses are not logged verbatim, the path explosion problem still applies when attempting to determine which return in the chain was corrupted, preventing accurate root cause analysis by $\sf{\mathcal{V}rf}$.
The discussion above highlights that $\sf{\mathcal{V}rf}$’s ability to learn an attack’s behavior and its root cause varies. Put simply, E1 only allows $\sf{\mathcal{V}rf}$ to learn if the reported path is invalid. However, it is impossible to learn what the invalid control flow path was or how it was exploited.
With E2, an invalid path is determined when a forward-edge target does not match possible successors in the CFG or when a return address in E2 does not match the value popped from the SS. In either case, $\sf{\mathcal{V}rf}$ can determine the exact indirect call or return exploited since all of these destinations are included in $\mathcal{C}F_{Log}$.
In E3, since $\sf{\mathcal{V}rf}$ only receives a hash chain of all return addresses, it cannot determine which return was corrupted or its destination address after the corruption. As a consequence, $\sf{\mathcal{V}rf}$ cannot make sense of subsequent forward edges in $\mathcal{C}F_{Log}$ because it cannot verify that $\sf{\mathcal{P}rv}$ has followed the expected control flow transfers (as a previously exploited return could have gone anywhere). At this point, continuing to traverse the CFG yields no additional information. Thus, $\sf{\mathcal{V}rf}$ can not learn anything more about the execution, resulting in limited analysis capability beyond detecting that some control deviation occurred. In sum, similar to E1, E3 results in $\sf{\mathcal{V}rf}$ being unable to examine exploit root causes.
It follows from the discussion above that E2 is the most adequate to support auditing and exploit root cause analysis by $\sf{\mathcal{V}rf}$ based on $\mathcal{C}F_{Log}$. Based on this conclusion, Sec. 4 presents SABRE to showcase how $\sf{\mathcal{V}rf}$ analysis can be performed based on a $\mathcal{C}F_{Log}$ when it includes a verbatim sequence of control flow transfers.
4 SABRE
SABRE showcases $\mathcal{C}F_{Log}$-based root cause analysis and remediation with verbatim control flow traces and without the source-code of the attested program. Fig. 3 shows SABRE workflow consisting of five internal modules: Static Analysis, Path Verifier, Exploit Locator, Patch Generator, and Patch Validator. SABRE takes as input an application binary (App.elf) and a verbatim $\mathcal{C}F_{Log}$ (E2 evidence type). The notation used in the rest of the paper is summarized in Tab. I.
Table I: Notation Summary
Symbol Description
$\mathcal{C}F_{Log}$ Verbatim CFA evidence (type E2)
App.elf The binary of the attested application
CFG The Control Flow Graph of App.elf
SS Shadow Stack of return addresses
$loop_{count}$ a special entry in $\mathcal{C}F_{Log}$ to optimize static loops (i.e., loops with no internal branches) used by prior works [49, 6]; it denotes the total times the previous $\mathcal{C}F_{Log}$ entry repeated
$addr_{target}$ The control data (e.g., return address, indirect jump target, or function pointer) that was corrupted
$\mathcal{C}F_{Slice}$ A slice of $\mathcal{C}F_{Log}$ that starts where $addr_{target}$ was initialized and ends where the corrupted $addr_{target}$ was used.
$base$ The symbol (i.e., the base address or register) storing $addr_{target}$
$MemMap$ map between memory addresses and values for memory addresses that are in use during Symbolic Data Flow Analysis.
$RegMap$ map between registers and values for registers that are in use during Symbolic Data Flow Analysis.
$X$ special symbolic value assigned to $base$ during Symbolic Data Flow Analysis to monitor memory accesses to it.
$addr_{Acc}$ the address of the memory instruction used to corrupt $addr_{target}$
$reg_{Acc}$ the register $addr_{Acc}$ dereferences to corrupt $addr_{target}$
$reg_{Acc}^{Init}$ the initial value of $reg_{Acc}$ before it is modified to dereference/corrupt $addr_{target}$. In the context of buffer overflows, it holds the start address of the overflown buffer.
$addr_{lower}$ the instruction address that sets the lower bound of the buffer
$addr_{upper}$ the instruction address that sets the upper bound of the buffer
4.1 Static Analysis
The first module of SABRE performs static analysis of the application binary (App.elf) to produce the CFG, which is used as input for subsequent modules. The CFG comprises nodes, where each node represents a set of contiguous (non-branching) instructions starting after the previous branch instruction and ending with the next branch, as shown in Fig. 4. This module is executed statically, once per App.elf. The remaining modules are executed for each received $\mathcal{C}F_{Log}$.
Figure 4: Example CFG and $\mathcal{C}F_{Log}$
4.2 Path Verifier
SABRE interprets each $\mathcal{C}F_{Log}$ entry as either: (1) a branch destination (denoted $addr_{Ni}$ in Fig. 4) or (2) a loop counter (denoted $loop_{count}$ in Fig. 4), indicating the number of times that a CFG node in the previous entry has executed. After receiving $\mathcal{C}F_{Log}$, Path Verifier uses this information to determine whether $\mathcal{C}F_{Log}$ constitutes a valid control flow path by traversing the CFG and SS, as described in Sec. 3.2. Upon detecting an invalid transfer, it proceeds to analyze the type of control flow instruction (i.e., ret or indirect call) whose destination address is being corrupted. This is done by first locating the node in the CFG that aligns with the last valid transfer in $\mathcal{C}F_{Log}$. The corrupted instruction is assigned as the last instruction of this CFG node. Using $App.elf$, the corrupted instruction and target address ($addr_{target}$) can be identified as either a return address or an indirect branch target. As a result, this module outputs the corrupted instruction and $addr_{target}$ to the next module of SABRE.
4.3 Exploit Locator
This module pinpoints exploit sources in three phases.
4.3.1 Phase 1. Backwards Traversal
Given the corrupted instruction from the Path Verifier, this phase attempts to find:
  1. (1)
    $\mathcal{C}F_{Slice}$: the section of $\mathcal{C}F_{Log}$ relevant to this particular exploit, which includes the $\mathcal{C}F_{Slice}$ entries between the time when memory storing $addr_{target}$ was assigned with the last pre-corruption value and the time that the corrupted branch executes causing the detected violation.
  2. (2)
    $base$: the memory address storing $addr_{target}$ at the beginning of $\mathcal{C}F_{Slice}$.
The procedure for determining $\mathcal{C}F_{Slice}$ and $base$ varies based on the type of $addr_{target}$. This is shown in Fig. 5 and Fig. 6 and described below:
Figure 5: Backward Traversal for a Return Address
Figure 6: Backward Traversal for an Indirect Call
  • $\bullet$
    Corrupted returns. Return addresses are initialized when pushed onto the stack during function calls, and they remain unmodified until the respective returns. Consequently, any corruption to a return address must have occurred between the function call and the return instruction. Therefore as a first step, App.elf is inspected to determine the start address of the function containing the corrupted return. In the example from Fig. 5, Path Verifier detects that the 17th $\mathcal{C}F_{Log}$ entry is invalid, corresponding to the corrupted return instruction at address 0xe152 and $addr_{target}$ pointing to address 0xf078. This module then inspects the disassembled instructions of App.elf and identifies that function F2 contains this corrupted instruction. Then, it saves the start address of F2, 0xe0b6, as the target of the backward traversal. A backward traversal of $\mathcal{C}F_{Log}$ is performed to find the last call to F2, which happens to be at the 8th $\mathcal{C}F_{Log}$ entry. As a result, this phase returns $\mathcal{C}F_{Slice}$ consisting of entries 8 to 17, i.e., $\mathcal{C}F_{Log}$[8:17]. In addition, since return addresses are always initialized on the stack, $base$ can be found in the stack pointer register (sp) at the beginning of $\mathcal{C}F_{Slice}$. Hence, this phase also outputs $base$ to be sp.
  • $\bullet$
    Corrupted forward edges. Indirect forward edge targets (indirect jumps/calls) are stored in a register operand at the time of their use. To identify $addr_{target}$’s initialization point in this case, $\mathcal{C}F_{Log}$ is used for a backward traversal of CFG while repeatedly following its assignments until reaching its initialization. For example, in Fig. 6, Path Verifier determines that, at instruction address 0xe150, $addr_{target}$ is stored in r15. Thus, this phase traverses instructions backward from address 0xe150 until it finds the latest assignment of r15 at address 0xe148. As r15 is assigned by -4(r4) at this address, Exploit Locator switches to tracking this new memory instead. This process continues until it encounters the instruction that initializes $addr_{target}$, which then sets the beginning of $\mathcal{C}F_{Slice}$ to the first $\mathcal{C}F_{Log}$ entry that includes this instruction. This phase finally assigns $base$ according to the type of the source operand in this instruction:
    1. (1)
[…]
    2. (2)
      a constant indicating a fixed memory address (i.e., when $addr_{target}$ is declared as a global variable). $base$ is set to this fixed address.
    3. (3)
      a return value from malloc (i.e, in the case where $addr_{target}$ is dynamically allocated). Backward traversal stops at the last malloc that defined it. Thus, $\mathcal{C}F_{Slice}$ only includes the malloc and free that potentially lead to the attack. In this case, $base$ is set to the return register from the malloc.
4.3.2 Phase 2. Symbolic Data Flow Analysis
Input : $\mathcal{C}F_{Slice}$: $\mathcal{C}F_{Log}$ slice containing vulnerability;
$App.elf$ : The application binary;
CFG: Control flow graph of $App.elf$;
$base$ : symbol holding $addr_{target}$ at $\mathcal{C}F_{Slice}$
start;
Output :  $addr_{Acc}$ : address of illegal memory instruction;
$MemMap$ : memory map containing
address-value pairs used in $\mathcal{C}F_{Slice}$;
$RegMap$ : register map containing
register-value pairs used in $\mathcal{C}F_{Slice}$;
$FreeList$ : list of freed pointers during
evaluation of $\mathcal{C}F_{Slice}$;
$MemMap,RegMap,FreeList\leftarrow\emptyset,\emptyset,\emptyset$
$MemMap\leftarrow\mathtt{AddConsts}(App.elf)$
if $base\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ Register$ then
[…]
else
[…]
end if
foreach $entry\in\mathcal{C}F_{Slice}$ do
      if $entry$ is not $loop_{count}$ then
             $dest_{addr}\coloneqq entry$
             $max_{iters}\coloneqq 1$
      else
             $max_{iters}\coloneqq entry$
       end if
      foreach $i\in\{1,2,...,max_{iters}\}$ do
             foreach $instr_{addr}\in CFG[dest_{addr}]$ do
                   $MemMap,RegMap,FreeList\leftarrow\mathtt{eval}(instr_{addr},MemMap,RegMap,FreeList)\;$
                  if $\exists(a,b)\in MemMap$ s.t. $a=X$ then
                         $addr_{Acc}\coloneqq instr_{addr}$
                         return: $(addr_{Acc},MemMap,RegMap,FreeList)$
                  else
                         continue
                   end if
             end foreach
       end foreach
end foreach
Algorithm 1 Symbolic Data Flow Analysis.
Within $\mathcal{C}F_{Slice}$ determined by Phase 1, Exploit Locator must locate the instruction that corrupts $addr_{target}$, typically requiring emulation of $\mathcal{C}F_{Slice}$ execution. Without access to run-time data prior to $\mathcal{C}F_{Slice}$ execution, Exploit Locator emulates this execution using symbolic analysis. Unlike traditional symbolic execution, we do not use symbolic analysis to explore different paths. Instead, we use it to reconstruct all possible data flows during $\mathcal{C}F_{Slice}$ execution [14, 15] and monitor for the change in data flows that result in corruption of $addr_{target}$.
Algorithm 1 details this phase. The symbolic data flow analysis takes $\mathcal{C}F_{Slice}$, base, App.elf, and CFG as inputs. As a first step, it initializes the execution state using MemMap, RegMap, and FreeList, where:
  • $\bullet$
    MemMap represents the set of memory states affected by $\mathcal{C}F_{Slice}$ execution, mapping memory addresses to their values. It is initially set to empty, except for addresses storing read-only data that can be determined from App.elf (e.g., data in .rodata segment of App.elf).
  • $\bullet$
    RegMap is a map of registers and their values for registers used in $\mathcal{C}F_{Slice}$ portion of App.elf execution. In the initial state, Reg is empty.
  • $\bullet$
    FreeList is a list of all pointers freed during the $\mathcal{C}F_{Slice}$ portion of App.elf execution. FreeList is also empty in its initial state.
The concrete value of base is only available during run-time and cannot be determined from $\mathcal{C}F_{Slice}$. Therefore, this phase initializes base with a special symbolic value ($X$) and updates the state (either MemMap or RegMap) in which base is stored accordingly. For instance, in Fig. 5 and 6, as base can be found in sp at the beginning of $\mathcal{C}F_{Slice}$, it sets $Reg$[sp] to $X$.
After initialization, this phase proceeds with symbolic execution of $\mathcal{C}F_{Slice}$ by iterating through each $entry$ in $\mathcal{C}F_{Slice}$. Assuming that the current $entry$ corresponds to a branch destination, it locates the CFG node containing this destination and extracts all instructions from it. Each instruction is then evaluated based on the current program state: if the instruction depends on memory or registers not currently present in $MemMap$ or $RegMap$, symbolic values (independent from $X$) are assigned to those memory/registers before executing it; otherwise, the instruction is simply executed symbolically. As a result of symbolic execution, the evaluation updates $MemMap$ and $RegMap$ to reflect the new execution state. Additionally, FreeList is updated whenever an evaluated instruction calls free(). When this is detected (by inspecting the instruction type and its destination address), the input parameter for the free() is first obtained from the previous instruction. Then, the input parameter and the call instruction’s address are added to the FreeList.
Next, Exploit Locator uses the evaluation results to determine whether a memory write to $base$ has occurred in the current evaluation. This is done by checking if a new value has been inserted to address $X$ in the updated $MemMap$. If so, it indicates that the current instruction introduces a value (i.e., writes) to memory at $base$, leading to $addr_{target}$’s corruption. Exploit Locator then outputs the address of this instruction as $addr_{Acc}$, along with $MemMap$, $RegMap$, and FreeList to the next phase.
When no write to $X$ is detected, it continues with symbolic execution over the remaining $\mathcal{C}F_{Slice}$ entries. We also note that this analysis may encounter a $\mathcal{C}F_{Slice}$ entry in the form of $loop_{count}$ (as depicted in Fig. 4). In this case, it repeats symbolic evaluation over the same CFG node for $loop_{count}$ times.
4.3.3 Phase 3. Detecting Exploit Type
In its final phase, the Exploit Locator attempts to determine whether the control flow attack resulted from a use-after-free or a buffer overflow, using the output from the previous phase.
Use-after-free. The FreeList can be inspected to determine whether a use-after-free has occurred. Since the previous phase stopped at the point where base was overwritten, the Exploit Locator checks the contents of FreeList to see if it contains base. If base is in FreeList and it was illegally modified after it was freed, it is likely that base was corrupted due to a use-after-free. In this case, Exploit Locator outputs the candidate attack root cause as a use-after-free.
Buffer Overflows. The Exploit Locator assumes that an attack exploits a buffer overflow when base is not in FreeList and when the CFG node containing $addr_{acc}$ has been executed multiple times in $\mathcal{C}F_{Slice}$ (implying that this CFG node is part of a loop that performs repeated memory writes, eventually overflowing into the bounds of $addr_{target}$). In this case, Exploit Locator outputs the attack root cause as a buffer overflow.
Figure 7: Backward traversal for determining buffer bounds.
Figure 8: Example patch for use-after-free vulnerabilities
4.4 Patch Generator
The goal of this module is to implement a binary patch that fixes the detected vulnerability. The exact steps depend on the detected root cause (either use-after-free or buffer overflow).
4.4.1 Patching Use-After-Free
Two root causes lead to a use-after-free: the free and the use along $\mathcal{C}F_{Slice}$. The use of a freed pointer typically occurs when a new object is allocated and when the same freed pointer (i.e., $addr_{target}$ in our case) is indirectly redefined through this new object. Without source-code semantics (such as variables, structs, pointer types/names), it is challenging to determine whether altering the use of $addr_{target}$ will lead to unwanted changes to a program’s functionality. Therefore, SABRE targets the unintended free along slice for patching.
SABRE patches the use-after-free by selectively removing the offending free from App.elf by replacing it with nop instructions. An example of SABRE’s use-after-free patch is depicted in Fig. 8. In this example, call instruction at address e098 was identified in the previous step as the call to free() in $\mathcal{C}F_{Slice}$ that frees $addr_{target}$. Call instructions (e.g., call in MSP430 and bl in ARM) take a full address as a parameter and thus span two memory addresses. Therefore, the call is replaced with two nop instructions.
4.4.2 Patching Buffer Overflows
Generating a patch for buffer overflow vulnerabilities includes the following tasks:
  • T1:
    Estimate the upper and lower bounds of the buffer that overflows into $addr_{target}$.
  • T2:
    Rewrite the instructions in the original code to safely record the bounds.
  • T3:
    Copy the function containing $addr_{acc}$ to an empty region of memory, wrapping $addr_{acc}$ in new instructions that perform a bounds check.
  • T4:
    Replace the original call to the function from [T3] with a trampoline to the patched version.
Detecting Buffer Bounds
Buffer overflows are eliminated via bound checks before memory accesses according to the buffer size. From the previous phase, the Exploit Locator identified the offending instruction to be located at $addr_{Acc}$ and that it causes an overflow onto $base$. This suggests that a buffer exists next to $base$ that must have been overwritten by a loop before the overflow occurs. It also implies that the offending instruction overflowing to $base$ is the same as that used to write data to the buffer.
With these insights, backtracking to where the destination operand (storing the overflown address) is initialized should provide the start address of the overflown buffer. Using this address, valid buffer bounds can be estimated to insert bound checks to the binary accordingly. This phase aims at finding these bounds before a patch can be generated and applied in the next modules.
The example of Fig. 7 depicts the two steps required to obtain the buffer’s bounds. The previous phase has detected that the instruction mov r14, 0(r15) triggers a write to base, setting $addr_{Acc}$ to 0xe290. From this instruction, this phase knows that its destination register, r15, stores the write address pointing into the overflown buffer, and thus labels it as $reg_{Acc}$. To determine the start address of the overflown buffer, SABRE performs a backward traversal from 0xe290 until it finds the instruction that initializes the buffer’s lower bound.
As depicted in Fig. 7, the buffer’s definition is influenced by -4(r4) at 0xe27e, which is in turn initialized by r11 at 0xe206. Tracking of definitions continues in this manner until a known pointer value is reached, as described in Sec. 4.3.1. In this example, definitions are tracked until 0xe0e4 since the stack pointer (a known pointer value) has been reached. Thus, Exploit Locator saves 0xe0e4 as its initialization point, as $addr_{lower}$ (the address defining the lower bound of the exploited buffer) and saves the value of $reg_{Acc}$ (sp in this example) as the lower bound of the overflown buffer at this instruction.
After this step, SABRE looks forward in $\mathcal{C}F_{Slice}$ to find an instruction that establishes an upper bound for $reg_{Acc}$ used at $addr_{lower}$. For example, since $reg_{Acc}$ is the stack pointer in Fig. 7, SABRE stops at the next call instruction at address 0xe106, as the stack pointer value before this call will determine the upper bound of the current stack frame, providing an approximation of the buffer’s upper bound. Therefore in this example, the address 0xe104 (the instruction before the call) is saved as the location that defines the upper bound ($addr_{upper}$). Consequently, the value of $reg_{Acc}$ at $addr_{upper}$ is used as the upper bound. When $addr_{target}$ is defined within the same stack frame as $reg_{Acc}$, base is used as the buffer’s upper bound. Similarly, when $reg_{Acc}$ obtains its definition from another source (e.g., a fixed memory address or as an output from malloc), $addr_{upper}$ may not exist or cannot always be determined through this method. In this case, SABRE defaults to using base as the upper bound.
Although this strategy could overestimate (i.e., the estimated bounds may include other data objects in addition to the buffer of interest), narrowing the scope of the bounds any further cannot be done without source-code level semantics. For example, the instruction at 0xe0ea modifies the stack pointer by ten. However, it cannot be determined through the binary alone whether this is allocating space for a buffer of size ten or for all stack variables totaling ten memory addresses. Although this differentiation cannot be made through binary alone, SABRE’s approach guarantees overflows into $addr_{target}$ are blocked by limiting buffer overflows into the buffer’s own stack frame.
Recording the Bounds
After identifying $addr_{lower}$ and $addr_{upper}$, SABRE completes T1 and has estimated the bounds of the buffer that overflows into $addr_{target}$. To complete T2, the values of $reg_{Acc}$ must be saved at both locations. This is completed by placing a trampoline both at $addr_{lower}$ and $addr_{upper}$ to preserve their values in a dedicated register. Fig. 9 depicts the entire patch for buffer overflow vulnerabilities and shows how these instructions are added for recording the bounds.
Figure 9: Example patch for buffer overflow vulnerabilities.
In this example, the trampolines and new instructions to save the lower and upper bounds are depicted with the instructions highlighted in purple and orange, respectively. First, the original instructions at $addr_{lower}$ and $addr_{upper}$ are replaced with a trampoline into a new region in the binary. For the upper bound, placing the trampoline alters the alignment, thus the trampoline and a nop are inserted to replace two instructions. Once trampolined into the new region, the instructions that were replaced by the trampoline execute first. Then, $reg_{Acc}$ (sp in this example) is recorded into reserved registers r9 and r10 to save the lower and upper bounds, respectively. To ensure the bounds cannot be altered, SABRE does a pass over App.elf binary to replace any other uses of these registers to ensure they are reserved. After the reserved registers have been updated with $reg_{Acc}$ value, the new instructions trampoline back into the previous code.
Inserting the Bounds Check
Then, Patch Generator proceeds to patch the offending memory instruction by prepending it with instructions that check whether the address being accessed is outside the valid boundary. If it is, the offending instruction is skipped. For the patch depicted in Fig. 9, the prepended instructions are highlighted in red and correspond to:
  • $\bullet$
    cmp r15, r9; jc \$+12; A comparison between the memory address being accessed (via r15) and the lower bound of the buffer (preserved in r9). If r15 is less than r9, it will skip the memory write by jumping to an offset of +12.
  • $\bullet$
    cmp r15, r10; jnc \$+8; A similar comparison except to the upper bound of the buffer (preserved in r10). If r15 exceeds the upper bound, the memory write will be skipped by jumping to an offset of +8.
The example in Fig. 9 uses instructions compiled for TI-MSP430, but the same patch is applied by SABRE to ARM-Cortex M33 binaries by using equivalent instructions. For example, instead of using jc and jnc, the patch targeting an ARM-Cortex M33 binary uses bcc and bhi, respectively.
To complete T3, the entire function that contains the exploited memory instruction ($addr_{acc}$) is rewritten to include this bounds check. The modified “safe” version of this function is placed into a new region in unused memory. We chose to make a safe copy and do not alter the original function in place to avoid altering all uses of this function (as the vulnerability is specific to this usage and does not necessarily apply to all calls in the code). As such, the final task T4 ensures that the safe version of this function is called at the previous $addr_{target}$-corrupting call site $addr_{target}$.
Calling the Safe Function
To realize T4, Patch Generator rewrites the original call instruction that leads to the function containing the buffer overflow. It changes this call instruction to instead call the patched version. This is depicted in Fig. 9 by the instructions highlighted in green.
The call instruction at 0xe106 previously called to the “unsafe” function at 0xe2fc. In the patched version, this call instruction now calls 0xe61a, the location of the patched version of the function. As depicted, this new function now contains the bounds-checking logic from T3.
4.5 Patch Validator
Before deploying the patched software on $\sf{\mathcal{P}rv}$, it must be validated to ensure its effectiveness against the previously identified attack. To achieve this, this module first builds an updated CFG of the patched binary. Then, it remaps the destination addresses in $\mathcal{C}F_{Slice}$ to their respective addresses in the updated CFG. It also inserts new entries into $\mathcal{C}F_{Slice}$ as necessary to account for the patching behaviors (e.g., a trampoline to/from the patch). Patch Validator then performs symbolic data flow analysis over the translated $\mathcal{C}F_{Slice}$, similar to how it is done in Sec. 4.3. When the analysis reaches the end of $\mathcal{C}F_{Slice}$ without $addr_{target}$ being corrupted, Patch Validator considers this patch effective. When the corruption still occurs, the attack was likely caused by another vulnerability. In this case, SABRE generates a report indicating to $\sf{\mathcal{V}rf}$ that further manual analysis is required.
(a) Static Analysis: Build CFG
(b) Path Verifier: Verify $\mathcal{C}F_{Log}$
(c) Exploit Locator: Backwards Traversal
(d) Exploit Locator: Symbolic DF Analysis
(e) Patch Generator: Generate Patch
(f) Patch Generator: Update ELF
(g) Patch Validator: Translate $\mathcal{C}F_{Log}$
(h) Patch Validator: Build Patched CFG
(i) Patch Validator: Symbolic DF Analysis
Figure 10: SABRE submodule timing for detecting and patching crafted control flow attacks inserted into a set of BEEBs [67] MCU software. This figure compares detecting/patching times for buffer overflow (OVF) and use-after-free (UAF) vulnerabilities across two MCUs and CFA architectures: hardware-based CFA atop MSP430 [6] and TEE-based CFA atop ARM Cortex-M33 [10].
5 Implementation details
SABRE was implemented as a combination of Bash and Python and evaluated on an Ubuntu 18.04 machine with an Intel Core i7-4790 CPU at 3.60GHz and 32GB RAM. For the evaluation, we use applications from the BEEBs benchmark for embedded platforms [67]. These applications were compiled for MSP430 and ARM Cortex-M and executed atop unmodified CFA architectures, namely ACFA [6] and TRACES [10], to generate a dataset of executables and corresponding $\mathcal{C}F_{Log}$-s. We believe that SABRE concepts generalize to several CFA architectures and select ACFA and TRACES as case studies due to their open-source availability.
To obtain CFGs from executables, SABRE first runs objdump over the executables to decode the assembly instructions. Then, it builds the CFG from this disassembled file. SABRE uses SymPy [68] for evaluating symbolic memory addresses and Keystone assembler [69] for generating the ARM Cortex-M patch instructions. Since no MSP430 assembler exists in Python packages, we build upon and modify MSProbe [70] for the MSP430 patch instructions. Finally, SABRE uses components from the Python library ELFtools [71] to generate the final patched executable. SABRE proof of concept prototype is publicly available at [16].
Table II: Size of SABRE modules
SABRE Static Path Exploit Patch Patch Utilities Total
Module Analysis Verifier Locator Generator Validator
Lines of Code 790 221 1672 1073 941 523 9122
Size (KB) 34.9 8.55 69.5 42.2 39.7 19.2 215.1
6 Evaluation
6.1 Code Size and TCB
Tab. II presents the size of SABRE modules in lines of code and total file size for each SABRE module. Exploit Locator is the largest as it implements the backward traversal and symbolic data flow analysis, the most complex processes of SABRE. Since building a patched CFG is a part of Patch Validator, this support is included in its calculated size, making it the second largest.
6.2 Run-times of SABRE Sub-Modules
Next, we evaluate the time to run each module in SABRE after it has determined that received CFA evidence contains a malicious path. Given the lack of reproducible exploits and incomplete CVEs for MCUs software (e.g., missing source codes, broken links, etc.) [72], we craft control flow attacks that exploit buffer overflow and use-after-free vulnerabilities (e.g., following a similar methodology to recent work [12]). We then insert them into the example benchmark applications [67] to closely mimic existing MCU-based CVEs [48, 47, 73, 40, 39, 38, 74] to the best of our ability. First, the main function is inspected to determine all function calls. Then, we insert the vulnerability into the last function called by the main. This vulnerability can be used to overwrite $addr_{target}$ to point back to the main function, creating an infinite loop in the program. We then execute the vulnerable programs on two unmodified CFA architectures: a TEE-based approach built atop ARM Cortex-M33 (TRACES [10]) and a custom hardware-based approach built atop TI MSP430 (ACFA [6]). We then collect $\mathcal{C}F_{Log}$ from this execution and pass it to SABRE along with the binary of the modified application (App.elf). Upon receiving $\mathcal{C}F_{Log}$, SABRE executes its sub-modules for further analysis. Fig. 10 shows the run-time for each sub-module.
6.2.1 Static Analysis
The time required to build the CFG via static analysis depends on the total number of branch instructions. Building the CFG for the MSP430 binaries requires more execution time, albeit smaller executables, as shown in Fig. 10(a). This is due to jump destinations in disassembled MSP430 binaries being decoded with offsets, not full addresses [75]. Therefore, MSP430 binaries require additional processing steps to determine jump destinations from the offsets, unlike ARM binaries, which disassemble into full addresses [76]. Generating the CFG from MSP430 binaries required $\approx$66.27 to 70.21ms, whereas ARM binaries required $\approx$9.46 to 17.23ms, depending on the particular application.
6.2.2 Path Validator
The time to verify $\mathcal{C}F_{Log}$ is presented in Fig. 10(b). The verification time depends on the number of branch instructions executed by the program before the corrupted $addr_{target}$ is used. In most cases, the crafted OVF generates a larger $\mathcal{C}F_{Log}$ since it introduces more control flow transfers before corrupting $addr_{target}$ compared to the crafted UAF examples. Additionally, $\mathcal{C}F_{Log}$-s of ARM executables (generated by TRACES [10]) require longer run-time to verify than $\mathcal{C}F_{Log}$-s of MSP430 executables (generated by ACFA [6]). This is because $\mathcal{C}F_{Log}$-s from TRACES do not contain any static branch instruction destinations, whereas $\mathcal{C}F_{Log}$-s from ACFA contain all control flow transfers. Consequently, TRACES $\mathcal{C}F_{Log}$-s require more time to verify since more processing is required by $\sf{\mathcal{V}rf}$ to account for the missing entries, whereas ACFA $\mathcal{C}F_{Log}$-s can be naively followed by $\sf{\mathcal{V}rf}$ without additional effort. The detection time through $\mathcal{C}F_{Log}$ verification for the evaluated examples ranged from $\approx$1.14 to 34.74ms.
6.2.3 Exploit Locator
Fig. 10(c) and Fig. 10(d) show the run-time for the Exploit Locator main submodules: the Backwards Traversal and Symbolic Data Flow (DF) Analysis. The Symbolic DF Analysis over $\mathcal{C}F_{Slice}$ incurs the highest run-time from 0.868 to 19.5s. For the OVF example in ARM, $\mathcal{C}F_{Slice}$ of the jfdctint program covers the highest number of instructions (7,750 in total), and thus, it takes the longest time to execute. In MSP430, the aha-compress program (abbreviated compress) has the same characteristic.
Since emulation of memory and registers is not required for Backwards Traversal, its run-time is more efficient in comparison (ranging from $\approx$0.026 to 5.31ms). For a similar reason as described above, Backwards Traversal of UAF in jfdctint requires the longest run-time. The run-time of Backwards Traversal for cover also stands out. The cover application is designed to test the performance while traversing every path in multiple large switch statements. Thus, the run-time slowdown arises due to tracking definitions across the memory stores/loads of each switch case. This effect is especially apparent in ARM binaries, which implement more memory operations for switch statements than MSP430. Overall, the run-time of the Exploit Locator is mainly dominated by the Symbolic DF Analysis.
6.2.4 Patch Generator
Patches are generated by first constructing a patched version of the instructions pertaining to the vulnerability. Then, the ELF executable is updated by adding a new dedicated section with the generated patch code. Times for these steps for MSP430 and ARM Cortex-M33 binaries are shown in Fig. 10(e) and Fig. 10(f), respectively.
Overall, the UAF patch is produced faster than the OVF patch due to its simplicity. Generating the patches for ARM Cortex-M33 requires more runtime due to its larger instruction set, differences in the compilation process (i.e., functions in ARM Cortex-M tend to include more loads/stores), and differences in the tooling used to generate the binary for the patch.
6.2.5 Patch Validator
Run-times of each step to verify MSP430 and ARM Cortex-M33 patches are shown in Fig. 10(g), Fig. 10(h), and Fig. 10(i).
The time to generate the translated $\mathcal{C}F_{Log}$ depends on the number of new branch instructions introduced as a part of the patch. Consequently, translating $\mathcal{C}F_{Log}$ for the UAF test cases requires negligible runtime since they introduce no new branch instructions (e.g., $<$1.0ms to perform this check, as shown in Fig. 10(g)).
The time to generate the patched CFG depends on the architecture, as described in Sec. 6.2.1. For the UAF case, the runtime is approximately the same since the patch effectively removed one CFG node. The OVF case requires more time to rewrite additional nodes. This difference is more pronounced for MSP430 due to the differences in decoding of jump instructions, as described in Sec. 6.2.1.
Similar to the Exploit Locator, the execution time of the Path Verifier is dominated by the Symbolic DF Analysis (ranging from 0.82s to 14.3s, as shown in Fig. 10(i)). Nonetheless, this analysis is performed faster in the Patch Validator because the patch causes the previously vulnerable node to exit more quickly. The bottleneck of the Symbolic DF Analysis comes from emulating memory access using MemMap. SABRE also requires an additional check for each memory write to determine whether the write corrupts $addr_{target}$ (recall Sec. 4.3). Future work could improve the efficiency of modeling memory accesses in SABRE.
6.3 Patch Sizes
Tab. III shows the increase in binary size (for MSP430 and ARM Cortex-M33) before and after the OVF patch for each application. The baseline (i.e., pre-patched) binary size is larger in ARM Cortex-M33 than in MSP430 due to several reasons. First, compared to the 16-bit MSP430 ISA, ARM Cortex-M is 32-bit and thus requires a larger instruction size. Second, the ARM Cortex-M33 binaries require additional instrumentation by the underlying CFA architecture TRACES [10] for secure generation of run-time evidence. On the other hand, the MSP430 applications are executed alongside a hardware-based CFA architecture (ACFA) that requires no instrumentation and thus incurs no size overhead. Since the baseline size of MSP430 binaries is smaller, their relative post-patched increase is larger (4.58-32.7%). In contrast, the ARM Cortex-M33 patch incurs a smaller relative increase (1.87-3.68%). Since the UAF patch simply replaces a call instruction in place with an equivalent number of nop-s, it incurs no increase in the binary size. However, it does incur an additional data memory overhead at runtime due to the missing free. The data memory overhead depends on the size of the object that is no longer freed.
Table III: Binary size comparison (bytes) for OVF patches
Buffer Overflow
MCU Platform MSP430 ARM Cortex-M33
Application Original Patched Increase (%) Original Patched Increase (%)
libbs 808 1072 32.7% 5212 5404 3.68%
lcdnum 872 1136 30.3% 5692 5880 3.3%
jfdctint 5764 6028 4.58% 7292 7480 2.58%
fibcall 792 1056 33.3% 5308 5496 3.54%
crc32 1236 1500 21.4% 5228 5416 3.6%
cover 3516 3780 7.51% 10268 10460 1.87%
compress 2940 3208 9.12% 6796 6988 2.83%
7 Related Work
7.1 Control Flow Attestation and Auditing
C-FLAT [4] proposed the concept of CFA by leveraging ARM TrustZone to detect and log control flow transfers. In C-FLAT, the binary is instrumented at branching instructions to trap execution in a TrustZone-protected “Secure World”. Once in the Secure World, the branching information is added to a hash chain. At the end of execution, C-FLAT’s hash chain produces a unique value representing the attested program’s control flow path. Inspired by C-FLAT, many follow-on techniques combine instrumentation and TEE support to improve CFA [53, 54, 50, 63].
Several subsequent hardware-based CFA approaches were proposed to remove the requirement of binary instrumentation by presenting custom hardware extensions for detecting control flow transfers [62, 51, 52]. In hardware-based CFA, the hardware extension includes dedicated modules for identifying branch instructions as they execute. The branch information is hashed on-the-fly by extending the MCU with a hardware-based hash engine. Although these techniques remove the requirement for instrumentation, the hardware cost of a fully hardware-based solution is too costly for MCUs.
Tiny-CFA [49] reduces the hardware cost significantly by executing an instrumented binary atop APEX [60], a low hardware-cost PoX architecture. ACFA [6] presents a hybrid (hardware/software co-design) that does not use hash engines, reducing the hardware cost for creating the evidence while still removing the requirement for instrumentation. ACFA’s custom hardware actively interrupts execution to transmit fixed-sized evidence slices to $\sf{\mathcal{V}rf}$. It is the first technique to enable run-time auditing by guaranteeing the eventual delivery of evidence.
The majority of existing work in this area focuses on the correctness and security of $\sf{\mathcal{P}rv}$ implementation to generate run-time evidence [4, 53, 54, 50, 63, 62, 51, 52, 49, 6, 58, 57]. They often mention that with authentic evidence, $\sf{\mathcal{V}rf}$ can detect the validity of the reported path by utilizing data structures computed prior to run-time (e.g., CFG+SS [53, 54, 50, 63, 6, 49] or hash-sets [62, 52, 4, 58]) or by emulating execution [51, 59]. However, as demonstrated by this work, all techniques except for the one utilizing verbatim evidence cannot infer the precise execution path on $\sf{\mathcal{P}rv}$ from their reported evidence. As such, they are not suitable for conducting root cause analysis after detecting the exploit on $\sf{\mathcal{P}rv}$.
In terms of $\sf{\mathcal{V}rf}$ point-of-view, only two existing studies [11, 12] propose alternative techniques for the $\sf{\mathcal{V}rf}$ to determine path validity under different scenarios. ZEKRA [11] considers a context in which $\sf{\mathcal{V}rf}$ might also be untrusted and aims to learn secrets about the $\sf{\mathcal{P}rv}$’s program through the run-time evidence. To address this issue, they propose a protocol that outsources the verification of evidence to a semi-dishonest worker. This semi-dishonest worker evaluates the reported path by passing it through an arithmetic circuit representing the CFG and generating a (zero-knowledge) proof that the verification result was evaluated on the most recent path and the correct CFG. Unlike SABRE, ZEKRA changes the system model of CFA by assuming $\sf{\mathcal{V}rf}$ might be untrusted. ZEKRA’s zero-knowledge procedures are only concerned with determining path validity, while SABRE aims to conduct further root cause analysis from the reported path.
RAGE [12] presents a technique for path verification that does not require constructing a complete CFG. RAGE trains a Graph Neural Network (GNN) on a set of $\sf{\mathcal{P}rv}$’s known program traces and validates run-time evidence using the trained GNN. Because of this, RAGE presents a method in which no knowledge of the code (source code or binary) is required during the run-time validation on $\sf{\mathcal{V}rf}$, unlike SABRE and other techniques. In contrast, the verification in SABRE and CFG-based techniques does not require prior execution contexts. Although SABRE requires constructing a partial CFG, we show that this is feasible for low-end embedded applications, as apparent in Fig. 10(a). More importantly, RAGE does not support a root cause analysis from the received evidence, which is the main goal of this work.
7.2 Symbolic Execution/Root Cause Analysis
The concept of symbolic execution was first introduced given the challenge of testing the reliability of programs [77]. Prior to its introduction, common techniques were to select a finite set of inputs to test the program’s correctness (i.e., fuzzing) or perform a formal analysis of a program to determine if it meets a particular specification for all inputs. A challenge for these approaches comes from determining the set of inputs or formal specifications for evaluating the program. Symbolic execution provides an alternative approach by executing the program on symbolic inputs and evaluating the input conditions that lead to a particular execution path. As a result, one symbolic execution can represent an infinite set of executions on concrete inputs. Since this original work [77], further research addresses its challenges and demonstrates its use to verify security features/detect bugs.
Various tools exist to provide a symbolic execution engine [78, 79, 80, 81], and their ability to symbolically explore program states has been leveraged in prior works across multiple platforms to evaluate their security. Some examples include bug testing/vulnerability detection [82, 83, 84, 85, 86, 87, 88], root cause analysis [14, 15], and verification of security properties [89, 90]. Since symbolic execution can be complex due to challenges with modeling memory, maintaining states of complex (or infinite) programs, or modeling behaviors of inputs/peripherals, Concolic Execution [91, 92, 93] combines characteristics of program emulation and symbolic execution by leveraging concrete values to guide symbolic path exploration.
Root Cause Analysis [94, 95, 14, 96, 97, 98] is an automated process for identifying the source of software bugs/unexpected behavior in software. ARCUS [14] is a root cause analysis framework with the goal of identifying the sources of run-time exploits in production systems. ARCUS contains separate modules to detect buffer overflows (stack & heap), integer over/underflows, use-after-free vulnerabilities, and string format vulnerabilities. It uses a hardware tracing module, Intel PT, to record a trace of control flow transfers whenever a program executes. The root cause analysis starts when the underlying CFI technique notifies ARCUS of a control flow violation. ARCUS then collects the execution trace from Intel PT and uses it to emulate the malicious execution. To ensure accurate emulation, a snapshot of the initial memory state before the malicious execution is required, and this snapshot is taken whenever a new thread to run an application program is created. While sharing the same goal of the root cause identification, SABRE operates on much more limited information (i.e., without a snapshot of initial memory as it is not available from $\mathcal{C}F_{Slice}$) as well as targets different classes of devices (i.e., Intel desktops in ARCUS vs ARM/MSP430 embedded devices in SABRE).
BENZENE [98] performs Root Cause Analysis starting from the application binary and crash-inducing inputs. It uses a data-flow graph (DFG) to track all values that were influenced by the inputs and mutates states along these paths to track when the program terminated successfully or crashed again. It then analyzes the crash and non-crash conditions to infer about crash conditions and generates a list of likely data-flow events that are the root causes. Unlike BENZENE, SABRE does not require data inputs and does not require constructing a DFG. Instead, SABRE combines CFG traversal with symbolic data-flow analysis to determine memory instructions that corrupt control data. Both SABRE and BENZENE create a report of a root cause. Yet, unlike BENZENE, SABRE generates and verifies a patch for the exploited root cause. Additionally, BENZENE assumes that data inputs were known, while SABRE relies on symbolic data-flow analysis to account for lacking data inputs.
8 Conclusion
We study if and how run-time evidence produced by CFA architectures can be used to identify corrupted branch targets, pinpoint exploited memory instructions, and generate/verify patches. We analyze and classify types of run-time evidence generated by current CFA architectures and conclude that verbatim evidence provides $\sf{\mathcal{V}rf}$ with the most flexibility for remote analysis. We propose SABRE to verify the path validity and to enable $\sf{\mathcal{V}rf}$ to locate/patch vulnerabilities using only the application binary and run-time evidence. We implement and evaluate the public prototype of SABRE [16] to demonstrate its effectiveness and compatibility with hardware-based and TEE-based CFA architectures. Our results indicate the feasibility of exploit detection and root cause analysis from CFA-generated evidence (when reported in a verbatim/lossless manner), even in cases where source code is unavailable. Importantly, the type of remote analysis facilitated by CFA evidence would be infeasible locally at $\sf{\mathcal{P}rv}$. We believe that this initial exploration and results further encourage (much needed) research on the verification side of run-time attestation protocols. Among possible paths for future work, we highlight (1) root cause analysis and patching assuming vulnerability types other than buffer overflows or use-after-free-s (e.g., integer overflows/errors, double free); and (2) the detection and remediation of stealthier out-of-order execution attacks such as data-oriented programming.
Appendix A Control Flow Hijacking Examples
Listings 1 and 2 show examples in which $\sf{\mathcal{A}dv}$ can corrupt the target of an indirect branch without modifying/injecting code by exploiting a stack-based buffer overflow (Listing 1) and a use-after-free vulnerability (Listing 2).
A.1 Control Flow Attack via Buffer Overflow
Consider the stack overflow example in Listing 1 from CWE-121 [99] in which the buffer hostname is overflowed by $\sf{\mathcal{A}dv}$-controlled input. In this example, only 64 bytes are allocated on the stack for the buffer hostname. Yet, the $\sf{\mathcal{A}dv}$-controlled input user_supplied_addr points to the hostname that will be copied into that buffer. Therefore, $\sf{\mathcal{A}dv}$ could point to a hostname larger than 64 bytes. As a result, the $\sf{\mathcal{A}dv}$-controlled values would overflow into other stack variables when the hostname is copied into the buffer at line 10. Since hostname is allocated onto the stack, $\sf{\mathcal{A}dv}$ can launch a control flow attack by corrupting other indirect branch targets (e.g., return addresses or function pointers) that are also stored on the stack.
A.2 Control Flow Attack via Use-After-Free
Consider the example in Listing 2 inspired by Wickman et al. [100]. In this example, the function handle_net_input contains a use-after-free vulnerability. This function first allocates a new instance of the Parser struct (parser) onto the heap, which contains a function pointer handler (as seen in lines 6-10). After setting the fields in parser, a command cmd is read from the network on line 21. When cmd is INVALID, it executes a free of the parser object on line 24 and reads a new command from the network on line 25. If this next command is PARSE, a new Array object array is allocated on line 32, which is the same size as a Parser object. Consequently, array->buf now overlaps with the previous parser->handler function pointer. On line 34, input is received from the network to set array->buf. After that, the freed pointer parser->handler is used on line 36. As a result, an arbitrary memory address, based on the value of array->buf, is called. When $\sf{\mathcal{A}dv}$ controls the input used to set array->buf, they can arbitrarily decide which section of the code to execute due to the use-after-free of parser->handler.
1void host_lookup(char *user_supplied_addr){
2 struct hostent *hp;
3 in_addr_t *addr;
4 char hostname[64];
5 in_addr_t inet_addr(const char *cp);
6
7 validate_addr_form(user_supplied_addr);
8 addr = inet_addr(user_supplied_addr);
9 hp = gethostbyaddr(addr, sizeof(struct in_addr), AF_INET);
10 strcpy(hostname, hp->h_name);
11}
Listing 1: Stack-based Buffer Overflow Example [99]
1typedef struct{
2 long used;
3 char buf[24];
4} Array; // 32-byte
5
6typedef struct { // 32-byte
7 long status;
8 void *start, *current;
9 int (*handler)(void *buf);
10} Parser;
11
12enum Command { INVALID, PARSE, ... };
13int net_parser(void *buf);
14
15int handle_net_input(int client_fd) {
16 // allocation
17 Parser *parser = (Parser *)malloc(32);
18 parser->handler = &net_parser;
19
20 enum Command cmd = INVALID;
21 read(client_fd, &cmd, sizeof(cmd));
22
23 if (cmd == INVALID) {
24 free(parser);
25 read(client_fd, &cmd, sizeof(cmd));
26 }
27 .
28 .
29 .
30 if (cmd == PARSE) {
31 // re-allocation
32 Array *array = (Array *)malloc(32);
33 // parser->handler() changes
34 read(client_fd, array->buf, 24);
35 // use-after-free
36 parser->handler(array->buf);
37 free(parser);
38 }
39}
Listing 2: Use-after-free for Control Flow Hijacking