### v3

# On the Verification of Control Flow Attestation Evidence

Abstract.
In run-time attestation schemes, including Control Flow Attestation (CFA) and Data Flow Attestation (DFA), a remote Verifier ($\sf{\mathcal{V}rf}$) requests a potentially compromised Prover device ($\sf{\mathcal{P}rv}$) to generate evidence of its execution control flow path (in CFA) and optionally execution data inputs (in DFA). Recent advances in this space also guarantee that $\sf{\mathcal{V}rf}$ eventually receives run-time evidence from $\sf{\mathcal{P}rv}$, even when $\sf{\mathcal{P}rv}$ is fully compromised. Reliable delivery, in theory, enables run-time auditing in addition to attestation, allowing $\sf{\mathcal{V}rf}$ to examine run-time compromise traces to pinpoint/remediate attack root causes. However, $\sf{\mathcal{V}rf}$’s perspective in this security service remains unexplored, with most prior work focusing on the secure generation of authentic run-time evidence on $\sf{\mathcal{P}rv}$.
[…]
journalyear: 2025copyright: acmlicensedconference: 18th ACM Conference on Security and Privacy in Wireless and Mobile Networks; June 30-July 3, 2025; Arlington, VA, USAbooktitle: 18th ACM Conference on Security and Privacy in Wireless and Mobile Networks (WiSec 2025), June 30-July 3, 2025, Arlington, VA, USAdoi: 10.1145/3734477.3734710isbn: 979-8-4007-1530-3/2025/06
1. Introduction
Embedded devices exist at the edge of larger systems, often serving as remotely operated sensors and on-demand actuators. Due to low energy and cost requirements, they are commonly implemented using low-end micro-controller units (MCUs) that lack advanced architectural security features (e.g., virtual memory or memory management units). Consequently, MCUs are more vulnerable and frequently targeted by cyber attacks (Nafees et al., 2023; Kayan et al., 2022).
Given their low budget for preventative security features, MCUs need inexpensive means to prove their software integrity to back-ends that rely on their services (e.g., a device owner or control center). Towards this goal, Remote Attestation (RA(Coker et al., 2011) has been proposed to give a resource-rich back-end – called a Verifier ($\sf{\mathcal{V}rf}$) – the ability to remotely assess the software state of a resource-constrained Prover MCU ($\sf{\mathcal{P}rv}$). In RA, $\sf{\mathcal{V}rf}$ challenges $\sf{\mathcal{P}rv}$ to return a cryptographic proof of $\sf{\mathcal{P}rv}$’s currently installed software binary. Only after receiving an authentic proof will $\sf{\mathcal{V}rf}$ trust that $\sf{\mathcal{P}rv}$ is installed with the correct software.
Although RA can effectively detect illegal code modifications, it cannot detect run-time attacks that do not modify code (Abera et al., 2016). For example, an adversary ($\sf{\mathcal{A}dv}$) could exploit a memory vulnerability (e.g., a buffer overflow (Cowan et al., 2000)) to corrupt control data, including return addresses, indirect jump targets, and function pointers. Consequently, $\sf{\mathcal{A}dv}$ can chain out-of-order instruction sub-sequences to execute (often Turing-complete) malicious behavior (Bletsch et al., 2011; Roemer et al., 2012). Since these attacks do not modify code, they are oblivious to RA.
To address this limitation, RA can be extended into Control Flow Attestation (CFA). In CFA, $\sf{\mathcal{V}rf}$ obtains cryptographic proof of both the installed software image and the exact control flow path followed in the most recent execution of the software (Ammar et al., 2024). This evidence and the respective proof are generated/authenticated by a root of trust (RoT) within $\sf{\mathcal{P}rv}$ that securely records control flow transfer destinations. Therefore, CFA evidence can, in principle, allow $\sf{\mathcal{V}rf}$ to detect both code modification and control-flow hijacking attacks. Recent architectures augment CFA to obtain run-time auditing (Caulfield and otehrs, 2023; Caulfield et al., 2024b) by guaranteeing the delivery of CFA evidence to $\sf{\mathcal{V}rf}$ even when $\sf{\mathcal{P}rv}$’s software is compromised. See Sec. 2 for details on CFA and run-time auditing RoTs.
We observe that prior work in this area (Ammar et al., 2024) has focused on secure design/implementation of $\sf{\mathcal{P}rv}$ RoTs, often overlooking $\sf{\mathcal{V}rf}$’s role. To our knowledge, no systematic analysis has explored $\sf{\mathcal{V}rf}$’s ability to interpret CFA evidence beyond basic path validation (i.e., determining whether the reported control-flow path is valid). However, in practice, after deeming a path invalid (thus $\sf{\mathcal{P}rv}$ compromise), $\sf{\mathcal{V}rf}$ would need to use CFA evidence to (1) identify the root cause of this compromise and (2) implement patches to mitigate it. Neither of these has been thoroughly investigated in the literature. The challenge increases when $\sf{\mathcal{P}rv}$’s software contains proprietary dependencies/third-party libraries with unavailable source code: a common occurrence in embedded system software chains.
Given the prevalence of memory-unsafe languages in embedded software and their various associated exploits ((2023), CISA), we argue that automated methods to identify unknown memory safety vulnerabilities from attested run-time evidence can result in timely detection of otherwise oblivious attacks and elimination of their root causes. While techniques have been proposed for root cause vulnerability analysis in high-end devices (e.g., based on Intel PT traces (Intel, 2015)), none of them can be used to analyze execution traces produced by CFA. In particular, these techniques (Yagemann et al., 2021a, b; Zeng et al., 2022; Liu et al., 2022) require initial memory snapshots for a remote server to perform a root cause analysis. This requirement is unlikely to be satisfied in the CFA settings for two reasons. First, this would require $\sf{\mathcal{P}rv}$ to transmit potentially large memory snapshots during CFA, conflicting with the resource-constrained nature of MCUs (Ammar et al., 2024). Second, this snapshot must be stored by $\sf{\mathcal{P}rv}$ during the attested execution. This, in turn, unnecessarily reduces memory available to regular tasks on $\sf{\mathcal{P}rv}$. We expand on these points in Sec. 6. Due to these conflicts, existing root cause analysis frameworks are not directly compatible with CFA. Our work bridges these gaps with the following contributions:
  1. (1)
    We analyze and classify the types of run-time evidence generated by existing CFA RoTs. Their trade-offs are examined in detail with particular focus on their ability to support $\sf{\mathcal{V}rf}$’s remote vulnerability detection and remediation.
  2. (2)
    As a case study, we present SABRE: a Security Analysis and Binary Repair Engine based on CFA evidence. SABRE combines binary analysis with CFA run-time evidence to automatically locate and patch buffer overflow and use-after-free vulnerabilities without requiring source-code knowledge.
  3. (3)
    Prior root cause analysis assumes $\sf{\mathcal{V}rf}$ has the entire execution context, including an initial memory snapshot. As discussed earlier, this assumption does not hold for CFA. To address this, SABRE takes a different approach: upon detecting corrupted control data (e.g., return addresses or indirect call targets), SABRE performs an additional backward definitions analysis to reconstruct a symbolic memory state leading to the corruption. Since the recovered state is not fully concrete, existing root cause analysis techniques cannot be directly applied. SABRE overcomes this by extending prior work’s forward symbolic data flow analysis to handle symbolic information. Based on this analysis and the detected vulnerability, SABRE automatically generates binary patches and validates their effectiveness.
  4. (4)
    We develop an open-source prototype of SABRE, supporting TI MSP430 and ARM Cortex-M MCU binaries (Caulfield et al., 2025) and run-time evidence generated by existing open-source run-time attestation RoTs (based on Trusted Execution Environments (TEEs) (Caulfield et al., 2024b) and custom hardware (Caulfield and otehrs, 2023)).
2. Background
2.1. Remote Attestation
RA is a challenge-response protocol between two parties: a remote $\sf{\mathcal{V}rf}$ and a potentially compromised $\sf{\mathcal{P}rv}$. It allows $\sf{\mathcal{V}rf}$ to remotely assess $\sf{\mathcal{P}rv}$ trustworthiness by measuring the content of $\sf{\mathcal{P}rv}$ program memory. As depicted in Fig. 1, a typical RA interaction involves the following steps:
  1. (1)
    $\sf{\mathcal{V}rf}$ requests attestation from $\sf{\mathcal{P}rv}$ by sending a cryptographic challenge $Chal$.
  2. (2)
    Upon receiving $Chal$, $\sf{\mathcal{P}rv}$ computes a measurement over its program memory and $Chal$ to produce report $H$.
  3. (3)
    $\sf{\mathcal{P}rv}$ sends the report $H$ back to $\sf{\mathcal{V}rf}$.
  4. (4)
    Upon receiving $H$, $\sf{\mathcal{V}rf}$ checks $H$ against the expected value to determine if $\sf{\mathcal{P}rv}$ has been compromised.
Early “software-based” attestation (Kennell and Jamieson, 2003; Seshadri et al., 2004, 2005, 2008) methods performed step (2) without relying on secrets and instead incorporated measurable architectural side-effects (e.g., execution/response time) along with the measurement of program memory. To make attestation more suitable to remote wireless connections with unpredictable network latency, subsequent methods required $\sf{\mathcal{P}rv}$ to respond with an authenticated integrity token — e.g., a message authentication code (MAC) or a digital signature – computed over program memory using a secret key (either a pre-shared symmetric key or the private key for a public key $\sf{\mathcal{V}rf}$ knows). Since threat models of RA (and of CFA consequently) assume that $\sf{\mathcal{P}rv}$ is susceptible to full software compromise, the secret key used in this operation must be securely stored by $\sf{\mathcal{P}rv}$’s RoT. Secure storage for the RA secret key implies some level of hardware support for the RoT implementation, ensuring that the key is unmodifiable and inaccessible to any untrusted software running on $\sf{\mathcal{P}rv}$. Therefore, “hardware-based” approaches incorporate dedicated hardware RoT for storing secret keys and computing the measurement (Petroni Jr et al., 2004; Kovah et al., 2012; Schellekens et al., 2008; Vliegen et al., 2019; Noorman et al., 2017). Since such hardware is too costly for low-end MCUs, recent approaches have proposed “hybrid” RoTs (Nunes et al., 2019; Eldefrawy et al., 2012; Brasser et al., 2015), which use hardware for storing secret keys and trusted software for performing the authenticated measurement.
2.2. Code Reuse and Control Flow Attacks
Control flow attacks are made possible due to an exposed memory vulnerability, such as common and pervasive instances like buffer overflow and use-after-free vulnerabilities ((2024), MITRE). Buffer overflows are common in embedded systems software since it is often written in unsafe memory languages like C/C++. For example, recent CVE’s (National Vulnerability Database, 2021b, 2020b, 2020c) have disclosed buffer overflows exploited for arbitrary code execution in IoT devices (Realtek Semiconductor Corp., 2021) and embedded real-time operating systems (Zephyr Project, 2016). Use-after-free is another class of vulnerabilities that can lead to full software compromise. As it is not uncommon for embedded software and RTOSes to use dynamic (i.e., heap) allocation (Masmano et al., 2003; Ramakrishna et al., 2008; Deligiannis and Kornaros, 2016), they can also be vulnerable to use-after-free attacks, which is evidenced by recently disclosed CVE’s (National Vulnerability Database, 2017, 2021a). Since these attacks do not require code modifications, they would remain undetected by classic RA mechanisms.
Verifier ($\sf{\mathcal{V}rf}$) (4) Verify Report (1) Request3) Report Prover ($\sf{\mathcal{P}rv}$) (2) Measure state
Figure 1. RA interaction
2.3. Run-time Attestation and Auditing
Run-time Attestation and Auditing are a set of security services that aim to generate and guarantee the delivery of accurate and authentic evidence of a $\sf{\mathcal{P}rv}$’s run-time behavior to a remote $\sf{\mathcal{V}rf}$, thus reflecting to $\sf{\mathcal{V}rf}$ whether any control flow attack has occurred. One type of run-time attestation is Control Flow Attestation (CFA(Nunes et al., 2021b; Sun et al., 2020; Dessouky et al., 2018; Zeitouni et al., 2017; Toffalini et al., 2019; Zhang et al., 2021; Abera et al., 2016; Geden and Rasmussen, 2019; Neto and Nunes, 2023; Yadav and Ganapathy, 2023; Armanuzzaman et al., 2025), which extends RA to detect run-time attacks in addition to illegal binary modifications. Upon receiving a request to execute some software function along with a RA challenge from $\sf{\mathcal{V}rf}$, $\sf{\mathcal{P}rv}$ executes the requested function, and a RoT in $\sf{\mathcal{P}rv}$ records a trace of this execution into protected memory. In the most fine-grained case, the trace is a control flow log ($\mathcal{C}F_{Log}$) of all control flow transfers that have occurred during the requested execution. After producing $\mathcal{C}F_{Log}$, $\sf{\mathcal{P}rv}$ performs the authenticated RA integrity measurement over the $\mathcal{C}F_{Log}$, the received challenge, and its own program memory to produce a timely and authenticated report. This report attests to the static state of $\sf{\mathcal{P}rv}$’s program binary and the run-time state within this timely execution. When applicable, data inputs – in Data Flow Attestation (DFA(Sun et al., 2020; Nunes et al., 2021a; Dessouky et al., 2018) – or produced outputs (Nunes et al., 2020; Caulfield et al., 2022) can also be included in this proof to bind obtained results to the proper execution of respective software functions on expected inputs.
Existing CFA techniques use either (1) binary instrumentation along with TEE support; or (2) custom hardware modifications to generate $\mathcal{C}F_{Log}$ by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that all branch instructions (e.g., jumps, returns, calls, etc.) are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends $\mathcal{C}F_{Log}$ to update the trace with the current branch destination. In hardware-based techniques, custom hardware interfaces with the MCU core to detect branches and records their destinations to a reserved memory region. Recent work (Caulfield and otehrs, 2023; Caulfield et al., 2024b) has also created mechanisms to implement reliable communication with $\sf{\mathcal{V}rf}$ as a part of the CFA RoT, enabling run-time auditing by ensuring evidence delivery even in the presence of malware that infects $\sf{\mathcal{P}rv}$ and refuses to deliver evidence about its actions/presence.
3. CFA Evidence and its Utility
3.1. Evidence Types
The various realizations of CFA RoTs (see Sec. 2) create and manage $\mathcal{C}F_{Log}$ differently, presenting trade-offs between the storage/communication cost and ease of verification/analysis by $\sf{\mathcal{V}rf}$. We classify existing methods in three types: hash-based evidence (E1), verbatim evidence (E2), and hybrid evidence (E3). Compared to Ammar et al. (Ammar et al., 2024), we further offer a more detailed analysis of each type’s suitability for advanced analysis.
Figure 2. A conditional branch (CB), indirect/direct calls (IC/DC), and return (RET) in (a) hash-based, (b) verbatim, and (c) hybrid evidence.
Hash-based Evidence (E1). As control flow transfers occur, branch destination addresses are accumulated into a single hash chain. This is performed by first accumulating the current destination address with the previous hash and then hashing the accumulated result (Abera et al., 2016; Dessouky et al., 2017; Zeitouni et al., 2017), as exemplified in Fig. 2.a. After the first branch occurs, its destination 0xe004 is accumulated into the hash. Since this is the first branch destination, it is accumulated with an initial constant to obtain $H_{0}$. Upon the second branch instruction, its destination address 0xf000 is accumulated with $H_{0}$ to obtain $H_{1}$. This process continues until the end of the requested execution generating a unique value that represents the path. Although simple control flows produce a single hash in E1, programs with more complex control flows (including those with nested loops and branching) might require a set of hash digests (Abera et al., 2016; Zeitouni et al., 2017). This approach yields minimal storage/transmission overheads (i.e., the fixed size of a hash digest). However, $\sf{\mathcal{V}rf}$ must perform its analysis solely based on the final hash digest and the executable binary.
Verbatim Evidence (E2). On the other side of the spectrum, more recent CFA techniques advocate for storing and transmitting $\mathcal{C}F_{Log}$ in its entirety (Nunes et al., 2021b; Dessouky et al., 2018; Toffalini et al., 2019; Caulfield and otehrs, 2023; Zhang et al., 2021; Caulfield et al., 2024b). As depicted in Fig. 2.b, the full sequence of destination addresses is maintained in a buffer. Therefore, this approach requires further (lossless) compression or is limited to small and self-contained operations to avoid filling the dedicated storage too quickly. A common strategy is to log simple loops (those without internal branches) with their address once and a count/condition pertaining to the number of iterations performed (Caulfield and otehrs, 2023; Nunes et al., 2021b; Zhang et al., 2021; Caulfield et al., 2024b). Another reduction is obtained by recording statically defined branch destinations as a single bit (Dessouky et al., 2018). Since static destinations can be determined from the binary alone, branch instructions with statically defined destinations can be recorded as a single bit (1/0) to determine if the branch was taken or not. Branches with dynamically defined destinations (e.g., returns and indirect calls) are still recorded with their full addresses. If $\mathcal{C}F_{Log}$ remains larger than available storage despite these optimizations, it can be split into a set of multiple fixed-size slices and streamed to $\sf{\mathcal{V}rf}$, as suggested in (Toffalini et al., 2019; Caulfield and otehrs, 2023; Caulfield et al., 2024b). Despite the aforementioned challenges, E2 provides $\sf{\mathcal{V}rf}$ with full control flow path information for analysis.
Hybrid Evidence (E3). Some techniques implement a combination of E1 and E2. OAT (Sun et al., 2020) and ARI (Wang et al., 2023) record forward edges verbatim and accumulate backward edges (returns) into a single hash chain. As shown in Fig. 2.c, the RoT in $\sf{\mathcal{P}rv}$ maintains a trace of forward edges ($T_{F}$) and a hash chain of returns ($H_{R}$). Since 0xf000 is the destination of an indirect call, its full address must be recorded whereas the first destination can be represented by a single bit ‘0’ encoding the ‘not-taken’ conditional branch address. After the attested execution ends, $T_{F}$ and $H_{R}$ are sent to $\sf{\mathcal{V}rf}$.
3.2. Detection & Remediation via E1, E2 & E3
Figure 3. SABRE sub-modules and internal workflow
Consider an attack detection strategy in which $\sf{\mathcal{V}rf}$ constructs the CFG of the executable binary, i.e., as suggested in (Abera et al., 2016; Caulfield and otehrs, 2023; Dessouky et al., 2017; Zeitouni et al., 2017; Toffalini et al., 2019; Zhang et al., 2021). An initial set of invalid paths is defined as forward edges that do not exist in the statically constructed CFG. As valid indirect targets (e.g., return addresses or indirect calls) are over-approximated using the static CFG (Burow et al., 2019), $\sf{\mathcal{V}rf}$ may also use $\mathcal{C}F_{Log}$ along with the program’s binary to emulate the execution of the program locally. In this process, a Shadow Stack (SS) can be emulated locally to verify backward edges (return destinations) and data flow (DF) analysis can be employed to determine a set of valid call sites for each indirect call, updating the CFG accordingly. We observe feasibility of this verification strategy varies greatly depending on evidence types E1, E2, and E3.
For E1, with the only evidence available being a hash digest, $\sf{\mathcal{V}rf}$ must infer which path has resulted in the digest. To that end, it must generate each possible legal path based on the CFG and use an emulated SS to accumulate the path’s hash during each step of the CFG traversal. However, the complexity of this task grows exponentially with the number of branches, leading to the intractable path explosion problem (Ramalingam, 1994; Baldoni et al., 2018). Considering anomalous/illegal paths, the analysis is even more challenging. $\sf{\mathcal{V}rf}$ would need to map the hash digest to the exact illegal path. Illegal paths do not exist in the CFG and thus must then be enumerated directly, making it effectively impossible to learn about the exploit behavior from this type of evidence (this precludes any analysis of exploit root causes, as discussed below).
Evidence types E2 and E3 allow $\sf{\mathcal{V}rf}$ to use reported traces to traverse the CFG, checking the reported destination against valid destinations. When there is a mismatch between the reported destination and the set of valid destinations for that branching instruction, $\sf{\mathcal{V}rf}$ determines that $\sf{\mathcal{P}rv}$ executed an invalid path. With both E2 and E3, $\sf{\mathcal{V}rf}$ is able to emulate an SS of return addresses. For verifying the integrity of returns in E2, $\sf{\mathcal{V}rf}$ checks each return destination individually as they are reached in the path traversal. The expected return address is popped from the SS and compared to the return address reported in $\mathcal{C}F_{Log}$, and an invalid path is detected when there is a mismatch. For E3, however, $\sf{\mathcal{V}rf}$ must recompute all returns reached into the final hash chain digest and compare it to the one reported in $\mathcal{C}F_{Log}$. If there is a mismatch of hashes, an invalid path is detected due to one of the return addresses. We note that verification based on E3 is only able to determine that some return address was corrupted. Since actual return addresses are not logged verbatim, the path explosion problem still applies when attempting to determine which return in the chain was corrupted, preventing accurate root cause analysis by $\sf{\mathcal{V}rf}$.
The discussion above highlights that $\sf{\mathcal{V}rf}$’s ability to learn an attack’s behavior and its root cause varies. Put simply, E1 only allows $\sf{\mathcal{V}rf}$ to learn if the reported path is invalid. However, it is impossible to learn what the invalid control flow path was or how it was exploited.
With E2, an invalid path is determined when a forward-edge target does not match possible successors in the CFG or when a return address in E2 does not match the value popped from the SS. In either case, $\sf{\mathcal{V}rf}$ can determine the exact indirect call or return exploited since all of these destinations are included in $\mathcal{C}F_{Log}$.
In E3, since $\sf{\mathcal{V}rf}$ only receives a hash chain of all return addresses, it cannot determine which return was corrupted or its destination address after the corruption. As a consequence, $\sf{\mathcal{V}rf}$ cannot make sense of subsequent forward edges in $\mathcal{C}F_{Log}$ because it cannot verify that $\sf{\mathcal{P}rv}$ has followed the expected control flow transfers (as a previously exploited return could have gone anywhere). At this point, continuing to traverse the CFG yields no additional information. Thus, $\sf{\mathcal{V}rf}$ can not learn anything more about the execution, resulting in limited analysis capability beyond detecting that some control deviation occurred. In sum, similar to E1, E3 results in $\sf{\mathcal{V}rf}$ being unable to examine exploit root causes.
It follows from the discussion above that E2 is the most adequate to support auditing and exploit root cause analysis by $\sf{\mathcal{V}rf}$ based on $\mathcal{C}F_{Log}$. Based on this conclusion, Sec. 4 presents SABRE to showcase how $\sf{\mathcal{V}rf}$ analysis can be performed based on a $\mathcal{C}F_{Log}$ when it includes a verbatim sequence of control flow transfers.
4. SABRE
SABRE showcases $\mathcal{C}F_{Log}$-based root cause analysis and remediation with verbatim control flow traces (E2) and without the source-code of the attested program. Fig. 3 shows SABRE workflow consisting of five internal modules: Static Analysis, Path Verifier, Exploit Locator, Patch Generator, and Patch Validator. SABRE takes as input an application binary (App.elf) and a verbatim $\mathcal{C}F_{Log}$ (E2 evidence type). The notation used in this work is summarized in Table 1.
4.1. Static Analysis and Path Verifier
The first two stages of SABRE are well studied from prior works (Ammar et al., 2024). The first module of SABRE performs static analysis of the application binary (App.elf) to produce the CFG used by subsequent modules. Each CFG node represents a set of contiguous (non-branching) instructions starting with the previous branch destination and ending with the next branch, as shown in Fig. 6. This module is executed once per App.elf. The remaining modules are executed for each received $\mathcal{C}F_{Log}$.
Figure 4. CFG and $\mathcal{C}F_{Log}$
Figure 5. Backward Traversal for a ret
Figure 6. Backward Traversal for an call
SABRE verifies the contents of the verbatim $\mathcal{C}F_{Log}$, the contents of which are assumed to have the format depicted in Fig. 6. SABRE interprets each $\mathcal{C}F_{Log}$ entry as either: (1) a branch destination (denoted $addr_{Ni}$ in Fig. 6) or (2) a loop counter (denoted $loop_{count}$ in Fig. 6), indicating the number of times that a CFG node in the previous entry has executed. After receiving $\mathcal{C}F_{Log}$, Path Verifier uses this information to determine whether $\mathcal{C}F_{Log}$ constitutes a valid control flow path by traversing the CFG and SS, as described in Sec. 3.2. Upon detecting an invalid transfer, it proceeds to analyze the type of control flow instruction (i.e., ret or indirect call) whose destination address is being corrupted. This is done by first locating the node in the CFG that aligns with the last valid transfer in $\mathcal{C}F_{Log}$. The corrupted instruction is assigned as the last instruction of this CFG node. Using $App.elf$, the corrupted instruction and target address ($addr_{target}$) are identified as either a return address or an indirect branch target. This module outputs the corrupted instruction and $addr_{target}$ to the next module of SABRE.
Table 1. Notation Summary
Symbol Description
$\mathcal{C}F_{Log}$ Verbatim CFA evidence (type E2)
App.elf The binary of the attested application
CFG The Control Flow Graph of App.elf
SS Shadow Stack of return addresses
$loop_{count}$ a special entry in $\mathcal{C}F_{Log}$ to optimize static loops (i.e., loops with no internal branches) (Nunes et al., 2021b; Caulfield and otehrs, 2023); denotes total times the previous $\mathcal{C}F_{Log}$ entry repeated
$addr_{target}$ The control data (e.g., return address, indirect jump/call target) that was corrupted
$\mathcal{C}F_{Slice}$ A $\mathcal{C}F_{Log}$ slice from $addr_{target}$’s initialization point its corrupted usage.
$base$ The symbol (i.e., the base address or register) storing $addr_{target}$
$MemMap$ address-value mapping for in-use memory addresses during Symbolic DF Analysis.
$RegMap$ register-value mapping for in-use registers during Symbolic DF Analysis.
$FreeList$ list of all freed pointers during Symbolic DF Analysis
$X$ special symbolic value assigned to $base$ during Symbolic DF Analysis.
$addr_{Acc}$ the address of the memory instruction used to corrupt $addr_{target}$
$reg_{Acc}$ the register $addr_{Acc}$ dereferences to corrupt $addr_{target}$
$reg_{Acc}^{Init}$ the initial value of $reg_{Acc}$ before it is modified to dereference/corrupt $addr_{target}$. In the context of buffer overflows, it holds the start address of the overflown buffer.
$addr_{lower}$ the instruction address that sets the lower bound of the buffer
$addr_{upper}$ the instruction address that sets the upper bound of the buffer
4.2. Exploit Locator
Phase 1. Backwards Traversal. Given the corrupted instruction from the Path Verifier, this phase attempts to find:
  1. (1)
    $\mathcal{C}F_{Slice}$: the section of $\mathcal{C}F_{Log}$ relevant to this particular exploit, including the $\mathcal{C}F_{Slice}$ entries between the time when memory storing $addr_{target}$ was assigned with the last pre-corruption value and the time that the corrupted branch executes causing the detected violation.
  2. (2)
    $base$: the memory address storing $addr_{target}$ at the beginning of $\mathcal{C}F_{Slice}$.
The procedure for determining $\mathcal{C}F_{Slice}$ and $base$ varies based on the type of $addr_{target}$. This is shown in Fig. 6 and Fig. 6.
Corrupted returns. Return addresses are initialized when pushed onto the stack during function calls, and they remain unmodified until the respective returns. Consequently, any corruption to a return address must have occurred between the function call and the return instruction. Therefore as a first step, App.elf is inspected to determine the start address of the function containing the corrupted return. In the example from Fig. 6, Path Verifier detects that the 17th $\mathcal{C}F_{Log}$ entry is invalid, corresponding to the corrupted return instruction at address 0xe152 and $addr_{target}$ pointing to 0xf078. This module then inspects the disassembled instructions of App.elf and identifies that function F2 contains this corrupted instruction. Then, it saves the start address of F2, 0xe0b6, as the target of the backward traversal. A backward traversal of $\mathcal{C}F_{Log}$ is performed to find the last call to F2, which happens to be at the 8th $\mathcal{C}F_{Log}$ entry. As a result, this phase returns $\mathcal{C}F_{Slice}$ consisting of entries 8 to 17, i.e., $\mathcal{C}F_{Log}$[8:17]. In addition, since return addresses are always initialized on the stack, $base$ can be found in the stack pointer register (sp) at the beginning of $\mathcal{C}F_{Slice}$. Hence, this phase also outputs $base$ to be sp.
Corrupted forward edges. Indirect forward edge targets (indirect jumps/calls) are stored in a register operand at the time of their use. To identify $addr_{target}$’s initialization point in this case, $\mathcal{C}F_{Log}$ is used for a backward traversal of CFG while repeatedly following its assignments until reaching its initialization. For example, in Fig. 6, Path Verifier determines that, at instruction address 0xe150, $addr_{target}$ is stored in r15. Thus, this phase traverses instructions backward from address 0xe150 until it finds the latest assignment of r15 at address 0xe148. As r15 is assigned by -4(r4) at this address, Exploit Locator switches to tracking this new memory instead. This process continues until it encounters the instruction that initializes $addr_{target}$, which then sets the beginning of $\mathcal{C}F_{Slice}$ to the first $\mathcal{C}F_{Log}$ entry that includes this instruction. This phase finally assigns $base$ according to the type of the source operand in this instruction:
  1. (1)
[…]
  2. (2)
    a constant indicating a fixed memory address (i.e., when $addr_{target}$ is declared as a global variable). $base$ is set to this fixed address.
  3. (3)
    a return value from malloc (i.e, in the case where $addr_{target}$ is dynamically allocated). Backward traversal stops at the last malloc that defined it. Thus, $\mathcal{C}F_{Slice}$ only includes the malloc and free that potentially lead to the attack. In this case, $base$ is set to the return register from the malloc.
Phase 2. Symbolic DF Analysis. Within $\mathcal{C}F_{Slice}$ determined by Phase 1, Exploit Locator must locate the instruction that corrupts $addr_{target}$, typically requiring emulation of $\mathcal{C}F_{Slice}$ execution. Without access to run-time data prior to $\mathcal{C}F_{Slice}$ execution, Exploit Locator emulates this execution using symbolic analysis. Unlike traditional symbolic execution (King, 1976), we do not use symbolic analysis to explore different paths. Instead, we use it to reconstruct all possible data flows during $\mathcal{C}F_{Slice}$ execution (Yagemann et al., 2021a, b) and monitor for the change in data flows that result in corruption of $addr_{target}$.
Algorithm 1 details this phase. The symbolic data flow analysis takes $\mathcal{C}F_{Slice}$, base, App.elf, and CFG as inputs. As a first step, it initializes the execution state using MemMap, RegMap, and FreeList, where:
  • $\bullet$
    MemMap represents the set of memory states affected by $\mathcal{C}F_{Slice}$ execution, mapping memory addresses to their values. It is initially set to empty, except for addresses storing read-only data that can be determined from App.elf (e.g., data in .rodata segment of App.elf).
  • $\bullet$
    RegMap is a map of registers and their values for registers used in $\mathcal{C}F_{Slice}$ portion of App.elf execution. In the initial state, Reg is empty.
  • $\bullet$
    FreeList is a list of all pointers freed during the $\mathcal{C}F_{Slice}$ portion of App.elf execution. FreeList is also empty in its initial state.
The concrete value of base is only available during run-time and cannot be determined from $\mathcal{C}F_{Slice}$. Therefore, this phase initializes base with a special symbolic value ($X$) and updates the state (MemMap or RegMap) in which base is stored accordingly. For instance, in Fig. 6 and 6, as base is found in sp at the start of $\mathcal{C}F_{Slice}$, $Reg$[sp] is set to $X$.
Input : $\mathcal{C}F_{Slice}$: $\mathcal{C}F_{Log}$ slice containing vulnerability;
$App.elf$ : The application binary;
CFG: Control flow graph of $App.elf$;
$base$ : symbol holding $addr_{target}$ at $\mathcal{C}F_{Slice}$ start;
Output :  $addr_{Acc}$ : address of illegal memory instruction;
$MemMap$ : memory map containing address-value pairs used in $\mathcal{C}F_{Slice}$;
$RegMap$ : register map containing register-value pairs used in $\mathcal{C}F_{Slice}$;
$FreeList$ : list of freed pointers during evaluation of $\mathcal{C}F_{Slice}$;
$MemMap,RegMap,FreeList\leftarrow\emptyset,\emptyset,\emptyset$
$MemMap\leftarrow\mathtt{AddConsts}(App.elf)$
if $base\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ Register$ then
[…]
else
[…]
end if
foreach $entry\in\mathcal{C}F_{Slice}$ do
      if $entry$ is not $loop_{count}$ then
             $dest_{addr}\coloneqq entry$
             $max_{iters}\coloneqq 1$
      else
             $max_{iters}\coloneqq entry$
       end if
      foreach $i\in\{1,2,...,max_{iters}\}$ do
             foreach $instr_{addr}\in CFG[dest_{addr}]$ do
                   $MemMap,RegMap,FreeList\leftarrow\mathtt{eval}(instr_{addr},MemMap,RegMap,FreeList)\;$
                  if $\exists(a,b)\in MemMap$ s.t. $a=X$ then
                         $addr_{Acc}\coloneqq instr_{addr}$
                         return: $(addr_{Acc},MemMap,RegMap,FreeList)$
                   end if
             end foreach
       end foreach
end foreach
Algorithm 1 Symbolic Data Flow Analysis.
After initialization, this phase proceeds with symbolic execution of $\mathcal{C}F_{Slice}$ by iterating through each $entry$ in $\mathcal{C}F_{Slice}$. Assuming that the current $entry$ corresponds to a branch destination, it locates the CFG node containing this destination and extracts all instructions from it. Each instruction is then evaluated based on the current program state: if the instruction depends on memory or registers not currently present in $MemMap$ or $RegMap$, symbolic values (independent from $X$) are assigned to those memory/registers before executing it; otherwise, the instruction is simply executed symbolically. As a result of symbolic execution, the evaluation updates $MemMap$ and $RegMap$ to reflect the new execution state. Additionally, FreeList is updated whenever an evaluated instruction calls free(). When this is detected (by inspecting the instruction type and its destination address), the input parameter for the free() is first obtained from the previous instruction. Then, the input parameter and the call instruction’s address are added to the FreeList.
Next, Exploit Locator uses the evaluation results to determine whether a memory write to $base$ has occurred in the current evaluation. This is done by checking if the value of $X$ in $MemMap$ has changed. If so, it indicates that the current instruction introduces a value (i.e., writes) to memory at $base$, leading to $addr_{target}$’s corruption. Exploit Locator then outputs the address of this instruction as $addr_{Acc}$, along with $MemMap$, $RegMap$, and FreeList to the next phase.
When no write to $X$ is detected, it continues with symbolic execution over the remaining $\mathcal{C}F_{Slice}$ entries. We note that this analysis may encounter a $\mathcal{C}F_{Slice}$ entry that is a $loop_{count}$ (as depicted in Fig. 6). In this case, it repeats symbolic evaluation on the same CFG node for $loop_{count}$ times.
Phase 3. Detecting Exploit Type. In its final phase, the Exploit Locator attempts to determine whether the control flow attack resulted from a use-after-free or a buffer overflow, using the output from the previous phase.
The FreeList can be inspected to determine whether a use-after-free has occurred. Since the previous phase stopped at the point where base was overwritten, the Exploit Locator checks the contents of FreeList to see if it contains base. If base is in FreeList and it was illegally modified after it was freed, it is likely that base was corrupted due to a use-after-free. In this case, Exploit Locator outputs the candidate attack root cause as a use-after-free.
The Exploit Locator assumes that an attack exploits a buffer overflow when base is not in FreeList and when the CFG node containing $addr_{acc}$ has been executed multiple times in $\mathcal{C}F_{Slice}$ (implying that this CFG node is part of a loop that performs repeated memory writes, eventually overflowing into $addr_{target}$). In this case, Exploit Locator outputs the attack root cause as a buffer overflow.
Figure 7. Example patch for use-after-free vulnerabilities
Figure 8. Backward traversal for determining buffer bounds.
4.3. Patch Generator
SABRE patching strategy (Duan et al., 2019; Xu et al., 2020; Eldefrawy et al., 2020; Zhang et al., 2017) redirects the vulnerable instructions to a secure implementation. The exact steps depend on the detected root cause (either use-after-free or buffer overflow).
4.3.1. Patching Use-After-Free
Two root causes lead to a use-after-free: the free and the use along $\mathcal{C}F_{Slice}$. The use of a freed pointer typically occurs when a new object is allocated and when the same freed pointer (i.e., $addr_{target}$ in our case) is indirectly redefined through this new object. Without source-code semantics (such as variables, structs, pointer types/names), it is challenging to determine whether altering the use of $addr_{target}$ will lead to unwanted changes to a program’s functionality. Therefore, SABRE targets the unintended free along slice for patching.
SABRE patches the use-after-free by selectively removing the offending free from App.elf by replacing it with nop instructions. An example of SABRE’s use-after-free patch is depicted in Fig. 8. In this example, call instruction at address e098 was identified in the previous step as the call to free() in $\mathcal{C}F_{Slice}$ that frees $addr_{target}$. Call instructions (e.g., call in MSP430 and bl in ARM) take a full address as a parameter and thus span two memory addresses. Therefore, the call is replaced with two nop instructions.
4.3.2. Patching Buffer Overflows
Generating a patch for buffer overflow vulnerabilities includes the following tasks:
  • T1:
    Estimate the upper and lower bounds of the buffer that overflows into $addr_{target}$.
  • T2:
    Rewrite the instructions in the original code to safely record the bounds.
  • T3:
    Copy the function containing $addr_{acc}$ to an empty region of memory, wrapping $addr_{acc}$ in new instructions that perform a bounds check.
  • T4:
    Replace the original call to the function from T3 with a trampoline to the patched version.
T1: Estimating Buffer Bounds. Buffer overflows are eliminated via bound checks before memory accesses according to the buffer size. From the previous phase, the Exploit Locator identified the offending instruction to be located at $addr_{Acc}$ and that it causes an overflow onto $base$. This suggests that a buffer exists next to $base$ that must have been overwritten by a loop before the overflow occurs. It also implies that the offending instruction overflowing to $base$ is the same as that used to write data to the buffer.
With these insights, backtracking to where the destination operand (storing the overflown address) is initialized should provide the start address of the overflown buffer. Using this address, valid buffer bounds can be estimated to insert bound checks to the binary accordingly. This phase aims at finding these bounds before a patch can be generated and applied in the next modules.
The example of Fig. 8 depicts the two steps required to obtain the buffer’s bounds. The previous phase has detected that the instruction mov r14, 0(r15) triggers a write to base, setting $addr_{Acc}$ to 0xe290. From this instruction, this phase knows that its destination register, r15, stores the write address pointing into the overflown buffer, and thus labels it as $reg_{Acc}$. To determine the start address of the overflown buffer, SABRE performs a backward traversal from 0xe290 until it finds the instruction that initializes its lower bound.
As depicted in Fig. 8, the buffer’s definition is influenced by -4(r4) at 0xe27e, which is in turn initialized by r11 at 0xe206. Tracking of definitions continues until a known pointer value is reached, as described in Sec. 4.2. In this example, definitions are tracked until 0xe0e4 since the stack pointer (a known pointer value) has been reached. Thus, Exploit Locator saves 0xe0e4 as $addr_{lower}$ (the address defining the lower bound of the exploited buffer) and saves the value of $reg_{Acc}$ (sp in this example) as the lower bound of the overflown buffer at this instruction.
After this step, SABRE looks forward in $\mathcal{C}F_{Slice}$ to find an instruction that establishes an upper bound for $reg_{Acc}$ used at $addr_{lower}$. For example, since $reg_{Acc}$ is the stack pointer in Fig. 8, SABRE stops at the next call instruction at address 0xe106, as the stack pointer value before this call will determine the upper bound of the current stack frame, providing an approximation of the buffer’s upper bound. Therefore in this example, the address 0xe104 (the instruction before the call) is saved as the location that defines the upper bound ($addr_{upper}$). Consequently, the value of $reg_{Acc}$ at $addr_{upper}$ is used as the upper bound. When $addr_{target}$ is defined within the same stack frame as $reg_{Acc}$, base is used as the buffer’s upper bound. Similarly, when $reg_{Acc}$ obtains its definition from another source (e.g., a fixed memory address or as an output from malloc), $addr_{upper}$ may not exist or cannot always be determined through this method. In this case, SABRE defaults to using base as the upper bound.
Although this strategy could overestimate (i.e., the estimated bounds may include other data objects and the buffer of interest), narrowing the bounds any further cannot be done without source-code level semantics. For example, the instruction at 0xe0ea modifies the stack pointer by ten. However, it cannot be determined through the binary alone whether this is allocating space for a buffer of size ten or for all stack variables totaling ten memory addresses. Although this differentiation cannot be made through binary alone, SABRE’s approach guarantees overflows into $addr_{target}$ are blocked by restricting overflows within the buffer’s stack frame.
T2: Recording the Bounds. After identifying $addr_{lower}$ and $addr_{upper}$, SABRE completes T1 and has estimated the bounds of the buffer that overflows into $addr_{target}$. To complete T2, the values of $reg_{Acc}$ must be saved at both locations. This is completed by placing a trampoline both at $addr_{lower}$ and $addr_{upper}$ to preserve their values in a dedicated register. Fig. 9 depicts the entire patch for buffer overflow vulnerabilities and shows how these instructions are added for recording the bounds.
In this example, the trampolines and new instructions to save the lower and upper bounds are depicted with the instructions highlighted in purple and orange, respectively. First, the original instructions at $addr_{lower}$ and $addr_{upper}$ are replaced with a trampoline into a new region in the binary. For the upper bound, placing the trampoline alters the alignment, thus the trampoline and a nop are inserted to replace two instructions. Once trampolined into the new region, the instructions that were replaced by the trampoline execute first. Then, $reg_{Acc}$ (sp in this example) is recorded into reserved registers r9 and r10 to save the lower and upper bounds, respectively. To ensure the bounds cannot be altered, SABRE does a pass over App.elf binary to replace any other uses of these registers to ensure they are reserved. After the reserved registers have been updated with $reg_{Acc}$ value, the new instructions trampoline back into the previous code.
Figure 9. Example patch for buffer overflow vulnerabilities.
T3: Inserting the Bounds Check Then, Patch Generator proceeds to patch the offending memory instruction by prepending it with instructions that check if the address being accessed is outside the valid boundary. If it is, the offending instruction is skipped. For the patch depicted in Fig. 9, the prepended instructions (highlighted in red) correspond to:
  • $\bullet$
    cmp r15, r9; jc \$+12; A comparison between the memory address being accessed (via r15) and the lower bound of the buffer (preserved in r9). If r15 is less than r9, it will skip the memory write by jumping to an offset of +12.
  • $\bullet$
    cmp r15, r10; jnc \$+8; A similar comparison except to the upper bound of the buffer (preserved in r10). If r15 exceeds the upper bound, the memory write will be skipped by jumping to an offset of +8.
The example in Fig. 9 uses instructions compiled for TI-MSP430, but the same patch is applied by SABRE to ARM-Cortex M33 binaries by using equivalent instructions. For example, instead of using jc and jnc, the patch targeting an ARM-Cortex M33 binary uses bcc and bhi, respectively.
To complete T3, the entire function that contains the exploited memory instruction ($addr_{acc}$) is rewritten to include this bounds check. The modified “safe” version of this function is placed into a new region in unused memory. We chose to make a safe copy and do not alter the original function in place to avoid altering all uses of this function (as the vulnerability is specific to this usage and does not necessarily apply to all calls in the code). As such, the final task T4 ensures that the safe version of this function is called at the previous $addr_{target}$-corrupting call site $addr_{target}$.
T4: Calling the Safe Function. To realize T4, Patch Generator rewrites the original call instruction that leads to the function containing the buffer overflow. It changes this call instruction to instead call the patched version. This is depicted in Fig. 9 by the instructions highlighted in green.
The call instruction at 0xe106 previously called to the “unsafe” function at 0xe2fc. In the patched version, this call instruction now calls 0xe61a, the location of the patched version of the function. As depicted, this new function contains bounds-checking from T3.
4.4. Patch Validator
Before deploying the patched software on $\sf{\mathcal{P}rv}$, it must be validated to ensure its effectiveness against the previously identified attack. To achieve this, this module first builds an updated CFG of the patched binary. Then, it remaps the destination addresses in $\mathcal{C}F_{Slice}$ to their respective addresses in the updated CFG. It also inserts new entries into $\mathcal{C}F_{Slice}$ as necessary to account for the patching behaviors (e.g., a trampoline to/from the patch). Patch Validator then performs symbolic DF analysis over the translated $\mathcal{C}F_{Slice}$, similar to how it is done in Sec. 4.2. When the analysis reaches the end of $\mathcal{C}F_{Slice}$ without $addr_{target}$ being corrupted, Patch Validator considers this patch effective. When the corruption still occurs, the attack was likely caused by another vulnerability. In this case, SABRE generates a report indicating to $\sf{\mathcal{V}rf}$ that further manual analysis is required.
libbslcdnumjfdctintfibcallcrc_32covercompress0020202020404040406060606080808080
(a) Build CFG
libbslcdnumjfdctintfibcallcrc_32covercompress0010101010202020203030303040404040
(b) Verify $\mathcal{C}F_{Log}$
libbslcdnumjfdctintfibcallcrc_32covercompress00222244446666
(c) Backwards Traversal
libbslcdnumjfdctintfibcallcrc_32covercompress05K10K15K20K
(d) Symb. DF Analysis
libbslcdnumjfdctintfibcallcrc_32covercompress00222244446666888810101010
(e) Patch Generation
libbslcdnumjfdctintfibcallcrc_32covercompress00202020204040404060606060
(f) Update ELF
libbslcdnumjfdctintfibcallcrc_32covercompress00111122223333
(g) Translate $\mathcal{C}F_{Log}$
libbslcdnumjfdctintfibcallcrc_32covercompress0020202020404040406060606080808080100100100100
(h) Build Patched CFG
libbslcdnumjfdctintfibcallcrc_32covercompress04K8K12K16K
(i) Symb. Patch Validation
Legend: ARM-UAF: ARM-OVF: MSP-UAF: MSP-OVF:
Figure 10. Run-time of SABRE submodules (ms) to analyze evidence of buffer overflow (OVF) and use-after-free (UAF) vulnerabilities on two MCUs and CFA architectures: hardware-based CFA atop MSP430 (Caulfield and otehrs, 2023) and TEE-based CFA atop ARM Cortex-M33 (Caulfield et al., 2024b).
5. Implementation & Evaluation
SABRE was implemented as a combination of Bash and Python and evaluated on an Ubuntu 18.04 machine with an Intel Core i7-4790 CPU at 3.60GHz and 32GB RAM. BEEBs benchmark applications for embedded platforms (Pallister et al., 2013) were used for the evaluation. They were compiled for MSP430 and ARM Cortex-M33 and executed atop unmodified CFA architectures, ACFA (Caulfield and otehrs, 2023) and TRACES (Caulfield et al., 2024b), to generate a dataset of executables and corresponding $\mathcal{C}F_{Log}$-s. The MSP430 prototype was configured with 8 KB of application program memory, and the ARM Cortex-M33 prototype was configured with 256 KB of application program memory. We believe that SABRE concepts generalize to several CFA architectures and use ACFA and TRACES as case studies due to their open-source availability.
To obtain CFGs from executables, SABRE first runs objdump over the executables to decode the assembly instructions. Then, it builds the CFG from this disassembled file. SABRE uses SymPy (Team, 2024) for evaluating symbolic expressions and Keystone assembler (Keystone, 2024) for generating ARM Cortex-M instructions. Since no MSP430 assembler exists in Python packages, we build upon and modify MSProbe (Swiftloke, 2024) for MSP430 instructions. Finally, SABRE uses the Python library ELFtools (Foundation, 2024c) to generate the final patched executable. SABRE proof of concept prototype is publicly available at (Caulfield et al., 2025).
5.1. Run-times of SABRE Sub-Modules
We evaluate the run-time of each SABRE sub-module after a malicious path was detected. Given the lack of reproducible exploits and incomplete CVEs for MCUs software (e.g., missing source codes, broken links, etc.) (Tan et al., 2024), we craft control flow attacks that exploit buffer overflow and use-after-free vulnerabilities, following a similar methodology to recent related work (Chilese et al., 2024). We then insert them into the example benchmark applications (Pallister et al., 2013) to closely mimic existing CVEs related to control flow hijacking and ROP in MCUs (National Vulnerability Database, 2021a, 2017, 2020a, 2020b, 2020c, b, 2022) to the best of our ability. For consistency across applications, we insert the vulnerability into the last function called by the main. The vulnerability is used to overwrite $addr_{target}$ to return/jump to the main function, causing an infinite loop. We execute the vulnerable programs on two unmodified CFA architectures – a TEE-based approach built atop ARM Cortex-M33 (TRACES (Caulfield et al., 2024b)) and a custom hardware-based approach built atop TI MSP430 (ACFA (Caulfield and otehrs, 2023)) – to collect $\mathcal{C}F_{Log}$-s of the attack, which are used by SABRE with the modified application binary (App.elf) for further analysis. Fig. 10 shows the run-time of each SABRE sub-module during the analysis.
Static Analysis. The time required to build the CFG depends on the number of branch instructions. Building the CFG for MSP430 binaries requires more run-time, albeit smaller executables, as shown in Fig. 10(a). This is due to jump destinations in disassembled MSP430 binaries being decoded with offsets, not full addresses (Foundation, 2024b). Therefore, MSP430 binaries require additional processing steps to determine addresses from the offsets, unlike ARM binaries, which disassemble into full addresses (Foundation, 2024a). Generating the CFG from MSP430 binaries required $\approx$66.27 to 70.21ms, whereas ARM binaries required $\approx$9.46 to 17.23ms, depending on the application.
Path Verifier. Fig. 10(b) presents the time to verify $\mathcal{C}F_{Log}$. The verification time depends on the number of branch instructions executed by the program before the corrupted $addr_{target}$ is used. In most cases, the crafted OVF generates a larger $\mathcal{C}F_{Log}$ since it introduces more control flow transfers before corrupting $addr_{target}$ compared to the crafted UAF examples. Additionally, $\mathcal{C}F_{Log}$-s of ARM executables (generated by TRACES (Caulfield et al., 2024b)) require longer run-time to verify than $\mathcal{C}F_{Log}$-s of MSP430 executables (generated by ACFA (Caulfield and otehrs, 2023)). This is because $\mathcal{C}F_{Log}$-s from TRACES do not contain any static branch instruction destinations, whereas $\mathcal{C}F_{Log}$-s from ACFA contain all control flow transfers. Consequently, TRACES $\mathcal{C}F_{Log}$-s require more time to verify since more processing is required by $\sf{\mathcal{V}rf}$ to account for the missing entries, whereas ACFA $\mathcal{C}F_{Log}$-s can be naively followed by $\sf{\mathcal{V}rf}$ without additional effort. The detection time through $\mathcal{C}F_{Log}$ verification for the evaluated examples ranged from $\approx$1.14 to 34.74ms.
Exploit Locator. Run-time for Backwards Traversal and Symbolic DF Analysis and are shown in Fig. 10(c) and Fig. 10(d), respectively. The Symbolic DF Analysis incurs the highest run-time from 0.868s to 19.5s. For the OVF example in ARM, $\mathcal{C}F_{Slice}$ of the jfdctint program covers the highest number of instructions (7750 in total), thus having the longest run-time. In MSP430, the aha-compress program (abbreviated compress) has the same characteristic.
Since tracking of memory and registers is not required for Backwards Traversal, its run-time is more efficient in comparison ($\approx$0.026 to 5.31ms). For a similar reason as described above, Backwards Traversal of UAF in jfdctint requires the longest run-time. The cover application is designed to test the performance while traversing every path in multiple large switch statements. Thus, the run-time slowdown arises due to tracking definitions across the memory stores/loads of each switch case. This effect is especially apparent in ARM binaries, which implement more memory operations for switch statements than MSP430. Overall, the run-time of the Exploit Locator is mainly dominated by the Symbolic DF Analysis.
Patch Generator. Overall, the UAF patch is produced faster than the OVF patch due to its simplicity, as shown in Fig. 10(e) and Fig. 10(f). Generating ARM Cortex-M33 patches requires more run-time due to its larger instruction set, differences in the compilation process (i.e., functions in ARM Cortex-M tend to include more loads/stores), and differences in the compiler tools.
Patch Validator. The time to translate the $\mathcal{C}F_{Log}$, shown in Fig. 10(g), depends on the number of new branch instructions introduced as a part of the patch. Therefore, translating $\mathcal{C}F_{Log}$ for the UAF test cases requires negligible run-time (e.g., $<$1.0ms) since it introduces no new branch instructions, whereas the OVF test case incurs time due to new branches introduces with patch.
The time to generate the patched CFG depends on the architecture. For the UAF case, the run-time is approximately the same since the patch effectively removed one CFG node. The OVF case requires more time to rewrite additional nodes. This difference is more pronounced for MSP430 due to decoding of jump instructions, as previously described for Static Analysis.
Like the Exploit Locator, the run-time of the Patch Validator is dominated by the Symbolic DF Analysis (from 0.82s to 14.3s in Fig. 10(i)). Nonetheless, this analysis is performed faster in the Patch Validator because the patch causes the previously vulnerable node to exit more quickly. The bottleneck of the Symbolic DF Analysis comes from emulating memory access using MemMap. SABRE also requires an additional check for each memory write to determine whether the write corrupts $addr_{target}$ (recall Sec. 4.2). Future work could improve SABRE’s modeling of memory accesses.
5.2. Patch Sizes
We compare program memory size before and after the patch is installed. As a baseline, the MSP430 binaries consumed from 792 to 5764 bytes (9.90% to 72.1% of the program memory) and the ARM Cortex-M33 binaries consumed from 5.3 KB to 10.3 KB (0.31% to 2.25% of the program memory). The difference in baseline is due to the MSP430 prototype being much more resource-constrained than the ARM Cortex-M33 prototype (see Sec. 5).
After the OVF patch, the MSP430 binaries consumed 1056 to 6028 bytes (13.2% to 75.4% of the program memory) and ARM Cortex-M33 binaries consumed 5.4 KB to 10.4 KB (from 0.41% to 2.35% of the program memory). This represents an average increase of $\approx$264.5 bytes ($\approx$3.31%) and $\approx$189.7 bytes ($\approx$0.10%), respectively.
Since the UAF patch simply replaces a call instruction in place with an equivalent number of nop-s, it incurs no increase in the binary size. However, it incurs additional data memory overhead at run-time due to the missing free. The data memory overhead depends on the size of the object that is no longer freed.
6. Related Work
Control Flow Attestation. Current CFA approaches generate execution path evidence combining binary instrumentation with TEE support (Armanuzzaman et al., 2025; Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield et al., 2024b; Yadav and Ganapathy, 2023) or using custom hardware to detect/log control flow transfers (Dessouky et al., 2017, 2018; Zeitouni et al., 2017; Caulfield and otehrs, 2023). Most CFA literature focuses on securely generating correct evidence on a compromised $\sf{\mathcal{P}rv}$, mentioning that this evidence can be used with CFG+SS (Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield and otehrs, 2023; Nunes et al., 2021b), hash-sets (Dessouky et al., 2017; Zeitouni et al., 2017; Abera et al., 2016; Yadav and Ganapathy, 2023), or execution emulation (Dessouky et al., 2018; Nunes et al., 2021a) to perform verification. As discussed in this paper, techniques that create hash-based/hybrid evidence (E1 and E3 from Sec. 3) cannot always infer the precise malicious execution paths on $\sf{\mathcal{P}rv}$ and are not suitable for root cause analysis.
To our knowledge, only two prior efforts (Debes et al., 2023; Chilese et al., 2024) propose approaches for CFA verification. ZEKRA (Debes et al., 2023) proposes a CFA verification in zero-knowledge proofs to ensure that an untrusted $\sf{\mathcal{V}rf}$ does not learn secrets from $\mathcal{C}F_{Log}$. RAGE (Chilese et al., 2024) presents a method for path verification by training a Graph Neural Network (GNN) on $\mathcal{C}F_{Log}$-s, removing the need for a CFG. Importantly, neither ZEKRA nor RAGE supports root cause analysis from received evidence, which is SABRE’s main goal. Although SABRE requires constructing a partial CFG (unlike RAGE), we show this is feasible for MCU applications; see Fig. 10(a). Ammar et al. (Ammar et al., 2024) provide a systematization of run-time defenses and discuss CFA verification as an open challenge. Unlike their discussion on the challenges in verifying CFA evidence, our work analyzes CFA evidence formats and their suitability for advanced analysis in depth. We also show how $\sf{\mathcal{V}rf}$ can use E2 evidence to pinpoint and remediate root causes concretely.
Root Cause Analysis. Root Cause Analysis frameworks  (Yagemann et al., 2021a; Liu et al., 2022; Park et al., 2024; Zeng et al., 2022; Zhou et al., 2022; Yagemann et al., 2021b) automate the process for identifying sources of software bugs or unexpected behavior. ARCUS (Yagemann et al., 2021a) collects an Intel PT trace and an initial memory snapshot, which are used by an ARCUS server after a run-time violation to identify root cause vulnerabilities during execution emulation and to generate a report. Due to the cost of storing and transmitting large memory snapshots, ARCUS and similar proposals (Yagemann et al., 2021b; Zeng et al., 2022; Liu et al., 2022; Zhou et al., 2022) are unlikely to integrate feasibly with CFA on low-end MCUs. BENZENE (Park et al., 2024) starts with the application binary and crash-inducing inputs, and it constructs a data-flow graph (DFG) to backtrack from crashes to track state conditions that lead to crashes and non-crashes. It then analyzes these conditions to generate likely root cause data-flow events. Unlike ACRUS and BENZENE, SABRE performs root cause analysis without large memory snapshots, crash inputs, or DFGs. Instead, SABRE combines CFG traversal with symbolic data-flow analysis to identify exploited instructions, generate a binary patch for the identified root cause, and verify its effectiveness.
Binary Rewriting and Patching. Automated binary rewriting  (Wenzl et al., 2019) can be used for pre-deployment security measures (Zhang and Sekar, 2015; Zhang et al., 2013; Wang et al., 2015; Wartell et al., 2012; Kiaei et al., 2021) and for post-deployment patch generation (Duan et al., 2019; Eldefrawy et al., 2020; Zhang et al., 2017; Rajput et al., 2023; Xu et al., 2020; Williams-King et al., 2020; Duck et al., 2020). Like prior work in this space (Duan et al., 2019; Xu et al., 2020; Eldefrawy et al., 2020; Zhang et al., 2017), SABRE patch instructions redirect execution to a secure implementation. However, SABRE identifies the target through root-cause analysis instead of function matching or other binary analysis methods. Real-world deployments assume the target device has a secure software update module to install the rewritten binary. For SABRE, no additional assumptions are required since this could be implemented as the remediation function within the run-time auditing workflow (Caulfield and otehrs, 2023; Caulfield et al., 2024b). Deployment challenges, like vendor-specific limits placed on software updates or flash endurance levels, would be accounted for in implementing the remediation function, and thus they are outside the scope of prior works in binary rewriting and SABRE.
7. Conclusion
We study how CFA evidence can be used to identify corrupted branch targets, pinpoint exploited memory instructions, and generate/verify binary patches. We analyze and classify types of CFA evidence and conclude that verbatim evidence provides $\sf{\mathcal{V}rf}$ with the most flexibility for remote analysis. We introduce SABRE to verify path validity and to locate/patch vulnerabilities using only the application binary and CFA evidence. We implement and evaluate SABRE’s public prototype (Caulfield et al., 2025), which demonstrates its effectiveness atop hardware-based and TEE-based CFA architectures.
Acknowledgments. We thank the anonymous reviewers and shepherd for their guidance.
APPENDIX
Appendix A Alternative Prover Configurations
Alternative mechanisms such as Address Space Layout Randomization (ASLR) (Shacham et al., 2004) or CFA optimizations (e.g., SpecCFA (Caulfield et al., 2024a)) may be concurrently employed on $\sf{\mathcal{P}rv}$ for increased security/performance. In both cases, SABRE remains effective as long as $\sf{\mathcal{V}rf}$ knows the configuration of the respective mechanism. For example, for CFA in MCUs (with no MMU), ASLR would be performed before $\sf{\mathcal{P}rv}$ deployment. Therefore, $\sf{\mathcal{V}rf}$ would have access to the randomized binary. Since an ASLR bypass would corrupt a pointer to a memory address post-randomization, SABRE would not need to perform any additional handling and can pass the randomized binary with the corresponding E2 evidence to SABRE. CFA optimizations on E2 evidence, including speculation-based ones such as SpecCFA (Caulfield et al., 2024a), are lossless by design. In it, $\sf{\mathcal{V}rf}$ replaces expected subpaths with reduced-sized reserved symbols. Since the subpath-to-symbol correspondence is known by $\sf{\mathcal{V}rf}$, its ability to analyze received evidence is not lost, regardless of this optimization.
Appendix B Limitations and Future Work
Scaling to Complex Systems. Most current CFA and run-time auditing targets MCUs (Ammar et al., 2024). Conversely, SABRE considers MCU binaries (i.e., MSP430 and ARM Cortex-M33). An open problem in this space is scaling $\sf{\mathcal{P}rv}$ RoTs from low-end MCUs to complex systems. Assuming CFA and generation of E2 in complex systems is possible, a scalability challenge would come from performing SABRE’s symbolic data-flow analysis. Although SABRE does not face typical path explosion risk in symbolic execution due to its use of concrete run-time traces to guide the execution, a unique concern would arise when emulating the active state (memory and registers) during symbolic data flow analysis, which for MCU binaries is already the bottleneck in SABRE’s Exploit Locator (recall Sec. 5.1). Optimizing SABRE’s state modeling during symbolic data-flow analysis is an interesting avenue for future work.
Adding Vulnerabilities and Attacks. This work focuses on control flow hijacking and code-reuse attacks. For this reason, we tailored our framework towards the most prevalent root-cause memory vulnerabilities for these attacks: use-after-free vulnerabilities and buffer overflow-induced out-of-bounds writes ((2024), MITRE). Additionally, these memory vulnerabilities are prevalent for launching arbitrary code execution in real-world MCU software (as discussed in Sec. 2.2). However, other memory vulnerabilities (e.g., signedness errors, integer overflows, and double frees) could lead to arbitrary code execution. In these cases, SABRE would determine that a generated patch was ineffective and generate a report to provide a starting point for manual inspection (as discussed in Sec. 4.4). Similarly, SABRE does not support data-oriented attacks (e.g., direct-data-manipulation or data-oriented-programming (Cheng et al., 2019)) and would fail in the same manner in its attempts to generate a patch or pass E2 verification if the attack did not corrupt any control-data. These attacks could be detected through DFA run-time evidence that includes data inputs (recall Sec. 2.3). Future work could extend SABRE to detect/patch other memory vulnerabilities and to use data inputs from DFA to address data-oriented attacks.